### v2

# Query Expansion in the Age of Pre-trained and Large Language Models: A Comprehensive Survey

Abstract.
Modern information retrieval (IR) must reconcile short, ambiguous queries with increasingly diverse and dynamic corpora. Query expansion (QE) remains central to alleviating vocabulary mismatch, yet the design space has shifted with pre-trained and large language models (PLMs, LLMs). In this survey, we organize recent work along four complementary dimensions: the point of injection (implicit/embedding vs. selection-based explicit), grounding and interaction (from zero-grounding prompts to multi-round retrieve–expand loops), learning and alignment (SFT/PEFT/DPO), and knowledge-graph integration. A model-centric taxonomy is also outlined, spanning encoder-only, encoder–decoder, decoder-only, instruction-tuned, and domain or multilingual variants, with affordances for QE such as contextual disambiguation, controllable generation, and zero-shot or few-shot reasoning. Practice-oriented guidance specifies where neural QE helps most: first-stage retrieval, multi-query fusion, re-ranking, and retrieval-augmented generation (RAG). The survey compares traditional and neural QE across seven aspects and maps applications in web search, biomedicine, e-commerce, open-domain question answering/RAG, conversational and code search, and cross-lingual settings. The survey concludes with an agenda focused on reliable, safe, efficient, and adaptive QE, offering a principled blueprint for deploying and combining techniques under real-world constraints.
Query Expansion, Information Retrieval, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Information retrieval query processing
1. Introduction
Modern IR systems must interpret increasingly short, context-poor queries issued via mobile, voice, and conversational interfaces, while the available content grows in scale, diversity, and linguistic complexity. The resulting vocabulary mismatch—users express intents with general or ambiguous terms whereas relevant documents use domain-specific terminology, paraphrases, or emerging jargon—is a long-standing obstacle. Context is often omitted, further hindering accurate intent inference. QE addresses these challenges by enriching the initial query with semantically related, contextually appropriate material so as to increase overlap with relevant documents.
Classical QE spans pseudo-relevance feedback (PRF), thesaurus/ontology-based expansion, translation-based models, and log-driven reformulation (Rocchio Jr, 1971; Voorhees, 1994; Lavrenko and Croft, 2017). Early benchmarks (e.g., TREC, CLEF) standardized evaluations and demonstrated the utility of PRF and semantic resources. Yet traditional techniques typically rely on static associations or shallow co-occurrence statistics and struggle with short, ambiguous, or long-tail queries (Carpineto and Romano, 2012; Voorhees, 1994). Topic drift, limited domain coverage, and the precision–recall trade-off remain recurring pain points.
Recent advances in PLMs and LLMs—e.g., BERT/RoBERTa-style encoders for context-sensitive representations (Nogueira and Cho, 2019), encoder–decoders such as T5/BART for controlled generation, and decoder-only LLMs (e.g., GPT-3/4, PaLM, LLaMA-family) for zero-/few-shot reasoning—have opened a broader design space for QE. These models support implicit expansion in the embedding space (e.g., PRF with dense vectors), selection-based term filtering with contextual encoders, and generative expansion via pseudo-documents or structured rationales. For instance, Query2Doc synthesizes pseudo-documents and yields 3–15% gains on MS MARCO and TREC DL (Wang et al., 2023c). Instruction tuning and in-context learning further improve format adherence and zero-shot robustness (Zhang, 2024). Open challenges persist: factual grounding, domain transfer under data sparsity, precision control for generative outputs, and cost/latency constraints at deployment.
Query expansion, as a classic problem in the field of information retrieval, has been systematically reviewed by numerous scholars. Early survey works, such as the study by Bhogal et al. (2007) provided a comprehensive overview of semantic QE; Sartori (2009) focused on ontology-driven QE, pointing out that ontologies can help obtain higher-quality expansion terms;Lei et al. (2011) summarizes the QE based on local analysis; Carpineto and Romano (2012) offered a brief summary of automatic QE, covering techniques based on linguistics, corpora, the query itself, retrieval logs, and Web data, and discussed related key issues; Rivas et al. (2014), confirmed that combining multiple classical techniques can effectively improve retrieval performance; Subsequently Raza et al. (2019) systematically discussed statistical QE including document analysis, retrieval and browsing log analysis, and Web knowledge analysis, while Azad and Deepak (2019) reviewed the core techniques, data sources, weighting and ranking methods, user involvement, and application scenarios of QE. However, these existing surveys primarily focus on traditional technical frameworks. In recent years, deep learning, represented by PLM and LLM, has achieved breakthroughs, bringing new opportunities and challenges to QE with their powerful semantic understanding and generation capabilities. Although new QE methods based on modern language models are emerging endlessly, we find that the academic community still lacks a survey that comprehensively covers this emerging technological wave. Therefore, this paper aims to fill this gap by systematically reviewing, classifying, and prospecting query expansion techniques that integrate modern PLM/LLM, providing a clear landscape and valuable reference for subsequent research in this field.
Scope and contributions
This survey synthesizes the landscape of QE in the PLM/LLM era with three unifying perspectives:
  1. (1)
    Paradigms. We trace the evolution from symbolic/statistical QE to implicit/embedding QE, selection-based explicit QE, and generative QE, highlighting how PLMs/LLMs reshape each paradigm.
  2. (2)
    Model taxonomy. We categorize PLMs/LLMs relevant to QE (encoder-only, encoder–decoder, decoder-only, instruction-tuned, domain/multilingual) and articulate their characteristic affordances for QE (disambiguation, controllable generation, zero-shot reasoning).
  3. (3)
    Practice and impact. We connect methods to deployment concerns (efficiency, latency, safety), compare traditional vs. neural QE, and review applications (web, biomedical, e-commerce, ODQA/RAG, conversational, code search).
Organization
Section 2 revisits foundations and traditional QE methods. Section 3 overviews pre-trained model families. Section 4 organizes Query Expansion techniques into four key dimensions: Point of Injection(implicit/embedding, selection-based explicit), Grounding and Interaction(integration of external knowledge and feedback), Learning and Alignment(training methods like SFT, PEFT, DPO), and KG-Augmented. Section 5 provides a comparative analysis of traditional vs. neural QE. Section 6 surveys application domains and use cases. Section 7 discusses open challenges and future directions.
[…]
Corpus-Wide Statistical Methods
Term Co-occurrence and Clustering:  Minker et al. (1972),  Willett (1988), Peat and Willett (1991), Global Similarity Thesaurus:  Qiu and Frei (1993)
Pre-defined Knowledge Resource
General-purpose Lexicons:  Voorhees (1993) Domain Ontologies:  Bhogal et al. (2007),  Lu et al. (2009),  Aronson and Rindflesch (1997), Fu et al. (2005) Local Analysis: Query-Dependent Evidence PRF/RF: Rocchio Jr (1971), Robertson and Jones (1976) LCA: Xu and Croft (2000, 2017) Query Logs and User Behavior Mining reformulations or Click-based semantic proximity:  Wen et al. (2002), Cui et al. (2002),  Cui et al. (2003),  Baeza-Yates et al. (2004),  Yin et al. (2009) Query Expansion of PLMs/LLMs Pre-trained Model Landscape Encoder-only: BERT (Devlin et al., 2019), RoBERT (Liu et al., 2019b) Encoder–Decoder: BART (Lewis et al., 2019), T5 (Raffel et al., 2019) Decoder-only LLMs GPT-3/4 (Floridi and Chiriatti, 2020; Achiam et al., 2023), PaLM (Anil et al., 2023), LLaMA-family Instruction-tuned Models: InstructGPT, FLAN-T5, LLaMA-2 Chat,(Ouyang et al., 2022; Wei et al., 2021; Touvron et al., 2023) Domain-specific / Multilingual: BioBERT(Lee et al., 2019), UMLS-BERT(Michalopoulos et al., 2020), SciBERT(Beltagy et al., 2019), multilingual BERT/RoBERTa families PLM/LLM-Driven QE Techniques
Point of injection
Implicit Embedding-based QE: ANCE-PRF (Yu et al., 2021), ColBERT-PRF (Wang et al., 2023b), Eclipse (D’Erasmo et al., 2025), QB-PRF (Zhang et al., 2024a), LLM-VPRF (Li et al., 2025c) Selection-based Explicit QE: CEQE (Naseri et al., 2021), SQET (Naseri et al., 2022), BERT-QE (Zheng et al., 2020), CQED (Khader and Ensan, 2023), PQEWC (Bassani et al., 2023)
Grounding and interaction
Zero-Grounding, Non-Interactive QE: LLM-Driven Single-Stage Expansion: Query2Doc (Wang et al., 2023c), CoT-QE (Jagerman et al., 2023), GAR (Mao et al., 2020), GRF (Mackie et al., 2023b), HyDE (Gao et al., 2023), Exp4Fuse (Liu and Zhang, 2025), Contextual clue sampling and fusion (Liu et al., 2022), SEAL (Bevilacqua et al., 2022) Grounding-Only, Non-Interactive QE: Corpus-Evidence Anchored Single-Pass Expansion: MILL (Jia et al., 2023), AGR (Chen et al., 2024), EAR (Chuang et al., 2023), GenPRF (Wang et al., 2023a), CSQE (Lei et al., 2024), MUGI (Zhang et al., 2024b), PromptPRF (Li et al., 2025b), FGQE (Jaenich et al., 2025) Grounding-Aware Interactive QE: Multi-Round Retrieve-Expand Loops: InteR (Feng et al., 2023), ProQE (Rashid et al., 2024), LameR (Shen et al., 2023), ThinkQE (Lei et al., 2025)
Learning and Alignment for QE with PLMs/LLMs
SoftQE (Pimpalkhute et al., 2024), RADCoT (Lee et al., 2024), ExpandR (Yao et al., 2025), AQE (Yang et al., 2025)
KG-augmented Query Expansion
Entity-based Expansion via Knowledge Graph Retrieval: KGQE (Perna, 2025), CL-KGQE (Rahman et al., 2019) Entity-based Expansion via Hybrid KG and Document Graph: KAR (Xia et al., 2025), QSKG (Mackie and Dalton, 2022) Application Domains Web Search Engines: EQE (Zhang et al., 2023) Biomedical Information Retrieval: Kelly (2021), Khader et al. (2022), Ateia and Kruschwitz (2025), Niu et al. (2024) E-commerce Search: BEQUE (Peng et al., 2024),  Wang et al. (2024b),  Zhang et al. (2022) Cross-Lingual and Multilingual Search: Rajaei et al. (2024) Open-Domain Question Answering: GAR (Mao et al., 2020), EAR (Chuang et al., 2023), AQE (Yang et al., 2025), MILL (Jia et al., 2023), AGR (Chen et al., 2024), HTGQE (Zhu et al., 2023), TDPR (Li et al., 2025a) Retrieval-Augmented Generation (RAG): QE-RAGpro (Hu et al., [n. d.]), QOQA (Koo et al., 2024), KG-Infused RAG (Wu et al., 2025),  Rajaei et al. (2024) Conversational Search: PRL (Mo et al., 2023), MISE (Kumar and Callan, 2020), QuReTeC (Voskarides et al., 2020) Code Search: GACR (Li et al., 2022), SSQR (Mao et al., 2023), SG-BERT/GPT2 (Liu, 2023), ECSQE (Bibi et al., 2025)
Figure 1. A Taxonomy of Query Expansion Techniques: From Traditional Methods to PLM/LLM-Driven Techniques and Applications.
2. Foundations of Query Expansion: Concepts and Traditional Methods
QE has been a central theme in classical IR (Bhogal et al., 2007; Sartori, 2009; Efthimiadis, 1996; Rivas et al., 2014; Raza et al., 2019; Lei et al., 2011; Carpineto and Romano, 2012; Azad and Deepak, 2019). Its goal is to increase lexical/semantic overlap between a user query and relevant documents, particularly when queries are short, ambiguous, or underspecified. Typical failures such as “car maintenance” missing “vehicle servicing” arise from vocabulary mismatch; QE mitigates this by adding related terms while preserving intent.
2.1. Concepts and Notation
Let the original query be a multiset of terms $Q$, stopwords $T^{\prime\prime}\!\subseteq\!Q$, and a set of candidate expansions $T^{\prime}$ from some source $D$. Following Azad and Deepak (2019), the expanded query is
(1) $Q_{\text{exp}}=(Q-T^{\prime\prime})\ \cup\ T^{\prime}.$
However, this set-based view is a simplification. In practice, QE is not merely about adding terms but also about re-weighting them to reflect their importance. The core difficulty is selecting $T^{\prime}$ that improves recall without distorting intent (Krovetz and Croft, 1992). Throughout, we write $D_{k}$ for top-$k$ feedback documents, and $\cos(\cdot,\cdot)$ is cosine similarity.
2.2. Global Analysis: Corpus-Wide and External Knowledge
Global methods operate on the assumption that stable, useful term associations can be inferred from large-scale evidence, independent of any single query. This evidence can be statistical patterns from the entire document corpus or curated knowledge from external resources.
2.2.1. Corpus-Wide Statistical Methods
This family assumes the target corpus carries sufficient regularities to infer useful associations.
Term Co-occurrence and Clustering
A classic approach infers related terms from co-occurrence statistics in documents. A common association is the cosine on binary/TF document incidence:
(2) $\mathrm{cos}(X,Y)=\frac{F(X,Y)}{\sqrt{F(X)\,F(Y)}},$
with $F(X)$ and $F(Y)$ denoting the document frequencies of terms $X$ and $Y$, respectively, and $F(X,Y)$ representing the co-occurrence count of $X$ and $Y$ in the same documents. While simple and unsupervised, Peat and Willett (1991) show a structural weakness: maximizing the cosine tends to select terms with similar (often high) document frequencies to the query term, which are weak discriminators under the Robertson–Sparck Jones (RSJ) theory (Robertson and Jones, 1976). Hence, co-occurrence expansion helps most for low-frequency query terms and can hurt for high-frequency ones.
Global Similarity Thesaurus
[…]
(3) $\mathrm{Sim}(q,t)=\sum_{t_{i}\in q}q_{i}\,S[t_{i},t].$
Selecting the top-$r$ terms yields a concept-consistent expansion list. Latent Semantic Analysis (LSA) (Abdelali et al., 2007)serves a similar purpose by mapping terms into a low-dimensional latent space, grouping semantically related terms together. These methods avoid hard clustering and allow variable-length expansions, but remain context-agnostic and can propagate popular-but-generic terms.
2.2.2. Methods Using Pre-defined Knowledge Resources
Curated resources inject explicit semantics and reduce ambiguity (Nasir et al., 2019).
General-purpose Lexicons (e.g., WordNet)
WordNet offers synsets and IS-A hierarchies for broad-coverage synonymy/hypernymy. However, short queries lack context for reliable WSD; naive synset expansion can underperform stemming/bag-of-words (Voorhees, 1993; Bhogal et al., 2007). Using glosses or intersecting sense neighborhoods can help (Navigli et al., 2003), but static coverage and sense ambiguity remain key issues for complex queries.
Domain Ontologies
Domain ontologies (e.g., UMLS/MeSH in biomedicine; legal/tourism ontologies) enable precise, concept-level expansion with fewer sense conflicts (Bhogal et al., 2007). Examples include mapping “heart attack” to “myocardial infarction” and related treatments/symptoms (Lu et al., 2009; Aronson and Rindflesch, 1997); composing tourism and geographic ontologies to resolve spatial language (Fu et al., 2005). Benefits are high precision and expert semantics; costs are curation/maintenance and domain dependence.
2.3. Local Analysis: Query-Dependent Evidence
In contrast to global methods, local analysis techniques derive expansion terms from a small set of documents retrieved in response to the current query. This approach is inherently context-aware but is highly dependent on the quality of the initial search results.
2.3.1. Relevance Feedback (RF) and Pseudo-Relevance Feedback (PRF)
RF adapts the query using assessed relevant/non-relevant documents; PRF automates this by assuming top-$k$ are relevant.
Vector-space (Rocchio)
RF relies on explicitly labeled document sets to refine the query. Given a set of relevant documents $D_{r}$ and an optional set of non-relevant documents $D_{nr}$, Rocchio (Rocchio Jr, 1971) proposes a vector-space update rule to construct the modified query vector:
(4) $\vec{q}_{m}=\alpha\vec{q}_{0}+\beta\frac{1}{|D_{r}|}\!\!\sum_{\vec{d}\in D_{r}}\!\vec{d}\;-\;\gamma\frac{1}{|D_{nr}|}\!\!\sum_{\vec{d}\in D_{nr}}\!\vec{d},$
where $\alpha,\beta,\gamma$ are non-negative weighting parameters: $\alpha$ controls the influence of the original query vector $\vec{q}_{0}$, $\beta$ adjusts the contribution of positive evidence from $D_{r}$, and $\gamma$ modulates the impact of negative evidence from $D_{nr}$. For PRF, only the positive centroid is used, with $D_{r}$ approximated by the top-$k$ initially retrieved documents $D_{k}$ as pseudo-relevant feedback.
Probabilistic (RSJ) weighting
Alternatively, expansion weights follow the RSJ discrimination:
(5) $w(t)=\log\frac{P(t\mid R)\,[1-P(t\mid NR)]}{P(t\mid NR)\,[1-P(t\mid R)]},$
prioritizing terms overrepresented in relevant documents (Robertson and Jones, 1976). In PRF, $P(t\!\mid\!R)$ is estimated on $D_{k}$ and $P(t\!\mid\!NR)$ against the background.
2.3.2. Local Context Analysis (LCA)
LCA refines PRF by requiring a candidate concept to co-occur with most terms in the original query ($Q$, with individual query terms $\{w_{i}\}$) within the local top-$n$ pseudo-relevant document set ($S$, i.e., the top-$n$ documents initially retrieved in PRF), reducing topic drift (Xu and Croft, 2000, 2017). Let “concepts” $c$ denote terms or phrases. We first define key metrics, where:
- $tf(x,y)$: Term Frequency of $x$ (concept/term) in document $y$, measured as the number of times $x$ appears in $y$ (length-normalized to avoid bias from long documents);
- $idf(x)$: Inverse Document Frequency of $x$, calculated as $\log_{10}\frac{N}{df(x)}$ (NN is total documents in the corpus, $df(x)$ is the number of documents containing $x$);
- $\delta$: A small positive smoothing parameter ($\delta>0$) to avoid zero values in product operations.
(6) $\operatorname{co}(c,w_{i})$ $=\sum_{d\in S}tf(c,d)\,tf(w_{i},d),$
(7) $\operatorname{co\_degree}(c,w_{i})$ $=\log_{10}(n)\,\log_{10}\!\big(\operatorname{co}(c,w_{i})+1\big)\cdot\mathrm{idf}(c),$
(8) $g(c,Q)$ $=\prod_{w_{i}\in Q}\big(\delta+\operatorname{co\_degree}(c,w_{i})\big),$
(9) $f(c,Q)$ $=\prod_{w_{i}\in Q}\big(\delta+\operatorname{co\_degree}(c,w_{i})\big)^{\mathrm{idf}(w_{i})},$
Concepts are ranked by $f(c,Q)$, and the top-$k$ concepts are appended to the original query for expansion. Empirically, the “soft-AND” product in $g(c,Q)$ and $f(c,Q)$ favors candidates supported by all query terms (rather than just individual terms), giving LCA stronger robustness against topic drift compared to frequency-only PRF methods.
2.4. Methods Using Query Logs and User Behavior
At web scale, logs capture implicit reformulations and intent signals (Wen et al., 2002; Cui et al., 2002, 2003; Baeza-Yates et al., 2004; Yin et al., 2009).
Mining reformulations
For a user’s original query $Q$, if a subsequent query reformulation $Q^{\prime}$ frequently follows $Q$ and leads to successful user interactions (e.g., clicks, long dwell time), terms $t$ that appear in $Q^{\prime}$ but not in $Q$ are treated as expansion candidates (Cui et al., 2002, 2003). To quantify the association between $Q$ and each candidate $t$, we use a statistic assoc(Q,t) that calculates the proportion of $Q$’s submissions where the subsequent $Q^{\prime}$ contains $t$:
(10) $\mathrm{assoc}(Q,t)=\frac{\mathrm{count}\big(Q\!\to\!Q^{\prime}\text{ with }t\!\in\!Q^{\prime}\big)}{\mathrm{count}(Q)}.$
Click-based semantic proximity
Queries leading to overlapping clicked document sets are semantically related. Let $D_{Q}$ denote the historical clicked set of a query $Q$ (i.e., all documents clicked by users after submitting $Q$); the semantic similarity between two queries $Q_{1}$ (to be expanded) and $Q_{2}$ is calculated as:
(11) $\mathrm{sim}(Q_{1},Q_{2})=\frac{|D_{Q_{1}}\cap D_{Q_{2}}|}{|D_{Q_{1}}\cup D_{Q_{2}}|},$
and terms from $Q_{2}$ (weighted by $\mathrm{sim}$ and within-$Q_{2}$ importance) can expand $Q_{1}$. These approaches are strongly data-driven but suffer from cold-start/long-tail sparsity and privacy constraints.
2.5. Limitations of Traditional QE
Although diverse in their approach, traditional QE techniques are bound by a shared, fundamental limitation: they operate on surface-level lexical statistics or static knowledge structures, lacking a deep model of dynamic, compositional semantics. This core deficiency manifests in several ways:
  1. (1)
    Context Insensitivity: Global methods, whether statistical or knowledge-based, are inherently context-agnostic. WordNet’s failure to resolve ambiguity in short queries is a direct consequence of this, as the correct ”sense” of a word is determined by its surrounding terms—a signal these models cannot effectively use.
  2. (2)
    Brittleness of Local Evidence: Local methods like PRF are highly susceptible to topic drift because they treat documents as bags of co-occurring words. They lack the ability to understand the compositional intent of a query like ”running shoes” and may drift towards the more general and popular topic of ”shoes” if the initial results are noisy.
  3. (3)
    Inability to Generalize: Log-based methods are powerful but cannot generalize to novel queries. They rely on observing past user behavior for specific query patterns and fail when encountering new combinations of terms that express a previously unseen intent.
Ultimately, traditional methods struggle to bridge the ”semantic gap” because they lack a true understanding of how language creates meaning through the combination and context of words. This persistent challenge set the stage for a paradigm shift in IR, creating a clear need for a new class of models that could learn deep, contextual, and compositional representations of language. These limitations motivate PLM/LLM-era methods that bring contextual encoding, generative reasoning, and stronger grounding to QE (see Sec. 4).
3. Pre-trained Model Landscape
Modern QE increasingly relies on PLMs and LLMs. While earlier neural encoders captured local semantics with limited context (Peters et al., 2018), contemporary pre-training paradigms deliver rich contextualization, generation, and instruction-following, which map naturally to key QE operations such as context-aware term selection, pseudo-document generation, and controllable reformulation. This section reviews the major PLM/LLM families and articulates their typical affordances for QE; concrete techniques are discussed in detail in Sec. 4.
3.1. Encoder-only
Encoder-only transformers (e.g., BERT, RoBERTa) are bidirectional masked LMs trained to produce context-sensitive token/sequence representations (Devlin et al., 2019; Liu et al., 2019b). For QE, their strengths are:
  • Context-sensitive disambiguation. For queries like “Apple launch event”, encoder-only models resolve sense and yield embeddings that align with Apple Inc. rather than the fruit, enabling safer selection-based expansion (e.g., synonym/phrase picking) instead of blind synonym injection.
  • Implicit/embedding-space expansion. They naturally support representation augmentation (e.g., PRF over dense vectors, token-level reweighting) without generating surface forms, which mitigates hallucination and reduces latency.
These models are a strong fit for PRF-style, selection-based, or reranking-centric QE where precision and efficiency are paramount (see Sec. 4).
3.2. Encoder–Decoder
Sequence-to-sequence (seq2seq) models pair a bidirectional encoder with an autoregressive decoder, combining understanding and generation. BART implements denoising autoencoding with span permutation/masking, synthesizing BERT-like encoding and GPT-like decoding (Lewis et al., 2019). T5 unifies tasks under a text-to-text objective and scales with large pre-training corpora (Raffel et al., 2019). For QE, encoder–decoders offer:
  • Direct generation of expansions. They produce answer-like pseudo-documents, titles, or rationales that enrich sparse/dense retrieval.
  • Structure-aware control. Prompts or templates can steer outputs to keywords, entities, or facets, improving interpretability and reducing topic drift.
Trade-offs include decoding latency and the need to control generation quality; however, they remain the dominant backbone for generative QE methods.
3.3. Decoder-only LLMs
Decoder-only transformers (e.g., GPT-3/4, PaLM, LLaMA-family) are trained autoregressively on massive corpora (Vaswani et al., 2017; Floridi and Chiriatti, 2020; Achiam et al., 2023; Anil et al., 2023). At sufficient scale, they exhibit emergent capabilities such as in-context learning and zero-shot reasoning (Wei et al., 2022). For QE, this translates to:
  • Zero-/few-shot expansion. Without task-specific fine-tuning, LLMs can infer latent intents and generate high-coverage reformulations or pseudo-documents.
  • Reasoning-guided reformulation. Chain-of-thought style prompts elicit structured expansions that surface latent entities/relations beneficial to retrieval.
Decoder-only LLMs offer the greatest flexibility but bring higher inference cost and require safeguards (e.g., grounding, self-checks) to prevent hallucinated entities in expansions.
3.4. Instruction-tuned & SFT/RLHF Models
Instruction-tuned models (e.g., InstructGPT, FLAN-T5, LLaMA-2 Chat) apply supervised fine-tuning (SFT) on instruction–response pairs and often reinforcement learning with human feedback (RLHF) for preference alignment (Ouyang et al., 2022; Wei et al., 2021; Touvron et al., 2023). Compared with base LMs, they:
  • Follow prompts reliably. Better adherence to requested output formats (keyword lists, entity tables, facet trees) is valuable for QE pipelines that fuse multiple signals.
  • Generalize in zero/few-shot settings. Aligned models often produce higher-precision expansions with fewer prompt heuristics (Wang et al., 2024a).
They are ideal front-ends for controlled generation (e.g., multi-query, rationale-then-keywords), especially when transparency and editability of expansions are required.
3.5. Domain-specific and Multilingual Models
Specialized pre-training narrows vocabulary mismatch in professional and non-English settings:
  • Domain-specific encoders. BioBERT adapts BERT on PubMed/PMC for biomedical text (Lee et al., 2019); UMLS-BERT injects UMLS CUIs/semantic types (Michalopoulos et al., 2020); SciBERT trains on scientific full text with SCIVOCAB to better cover scholarly terminology (Beltagy et al., 2019). These models consistently improve entity resolution and concept linking, which directly benefits QE in biomedical and scientific IR.
  • Multilingual encoders/LLMs. Multilingual variants (e.g., multilingual BERT/RoBERTa families) support cross-lingual QE by mapping semantically aligned terms across languages, enabling bilingual/multilingual reformulations without parallel supervision. In practice, they are used to (i) translate-and-expand or (ii) expand-then-translate within cross-lingual retrieval pipelines.
3.6. Architectural Affordances for QE: A Practitioner’s View
Table 1 summarizes typical affordances and trade-offs by model family. In short, encoder-only models excel at selection/embedding-space QE with strong efficiency; encoder–decoders add structured generation under tighter control; decoder-only LLMs unlock zero-shot, reasoning-driven expansion; instruction-tuned models improve format adherence and safety; domain/multilingual variants mitigate vocabulary and language gaps. Section 4 details concrete algorithms that instantiate these affordances.
Table 1. Pre-trained model families and typical QE affordances.
Family Pre-training gist QE strengths (typical use) Common trade-offs
Encoder-only (BERT,
RoBERTa) (Devlin et al., 2019; Liu et al., 2019b)
Bidirectional masked LM; strong contextual encodings Term/phrase selection, PRF in embedding space; robust disambiguation; efficient reranking Limited free-form generation;
depends on candidate sources
Encoder–Decoder
(BART, T5) (Lewis et al., 2019; Raffel et al., 2019)
Denoising/text-to-text seq2seq; encoder+decoder Structured generative QE:
pseudo-docs, titles, rationales;
promptable output control
Decoding latency;
quality control of generations
Decoder-only LLM
(GPT-3/4, PaLM, LLaMA)
(Floridi and Chiriatti, 2020; Achiam et al., 2023; Anil et al., 2023)
Autoregressive LM at scale; in-context learning (Wei et al., 2022) Zero-/few-shot expansion;
reasoning-guided reformulation;
broad domain coverage
Higher cost; hallucination risk without grounding
Instruction-tuned (InstructGPT, FLAN-T5, LLaMA-2 Chat) (Ouyang et al., 2022; Wei et al., 2021; Touvron et al., 2023) SFT + (optionally) RLHF;
preference-aligned outputs
Format adherence
(keywords/entities/facets); safer, more controllable expansions
May over-adhere to instructions; still needs retrieval grounding
Domain-specific / Multilingual (Lee et al., 2019; Michalopoulos et al., 2020; Beltagy et al., 2019) Further pre-training on target domains/languages Bridges domain/lingual vocabulary gaps; better entity/term recall Narrow transfer outside domain; requires domain corpora
Practical guidance
When precision and latency dominate (e.g., production web search, e-commerce), start with encoder-only selection/PRF and add small encoder–decoder components for targeted generation. For recall-limited or long-tail queries, leverage decoder-only LLMs or instruction-tuned seq2seq for zero-shot multi-query expansion, but incorporate corpus-steering/feedback (see Sec . 4) to curb drift. In domain or cross-lingual settings, prefer specialized encoders and multilingual models to reduce reliance on noisy translation.
From Model Selection to QE Implementation
The preceding guidance maps PLM/LLM families to retrieval priorities. Realizing their QE value requires addressing four key questions: how to inject expansion signals, ground expansions in corpus evidence, align generation with retrieval goals, and integrate structured knowledge. Chapter 4 operationalizes these as four axes to bridge model capabilities with practical techniques, forming a unified framework for neural query expansion. These questions and axes structure the technical discussion in Chapter 4.
4. PLM/LLM-Driven QE Techniques
Modern query expansion revisits a classic idea—add missing evidence to a query—under the lens of PLMs and LLMs. What changes in the LLM era is not the goal but where and how the additional signal is injected: into the query’s vector representation, into a curated set of explicit terms or spans, or into freshly produced text that stands in for what the user might have asked if they had been more specific. Equally important are the choices of grounding (does expansion come from the corpus or from a model’s world knowledge?), learning regime (prompt-only vs. alignment and distillation), and interaction pattern (single shot vs. feedback-in-the-loop).
Our organizing view
We structure the literature along four complementary axes that unify “traditional” QE with PLM/LLM-based methods.
  1. (i)
    Point of injection. Implicit, embedding-level methods strengthen the query vector or token embeddings without emitting new terms (e.g., ANCE-PRF, ColBERT-PRF, Eclipse, QB-PRF, LLM-VPRF). Selection-based explicit methods keep the vocabulary corpus-grounded by ranking terms or spans from pseudo-relevant documents or curated resources (e.g., CEQE, SQET, BERT-QE, CQED, PQEWC).
  2. (ii)
    Grounding and interaction. Zero-grounding, non-interactive approaches rely mainly on an LLM’s prior knowledge, sometimes with minimal corpus hints (e.g., Query2Doc, CoT-QE, HyDE, Exp4Fuse), and issue the expanded query once without additional retrieval–generation cycles. Grounding-only, non-interactive methods explicitly condition generation on pseudo-relevant documents or other corpus evidence in a single retrieve–generate–requery pass, often using constraints, selection, or feedback calibration to control drift (e.g., MILL, AGR, EAR, GenPRF, CSQE, MUGI, PromptPRF, FGQE). Grounding-aware interactive designs run multiple retrieve–expand loops (e.g., InteR, ProQE, ThinkQE), using feedback from earlier retrieval stages to guide later expansions.
  3. (iii)
    Learning and alignment. Beyond zero/few-shot prompting, systems align generation to retrieval utility via supervised fine-tuning, preference optimization, or distillation (e.g., SoftQE, RADCoT, ExpandR, AQE).
  4. (iv)
    Knowledge Graph augmentation. KG-augmented strategies are categorized into two types: those using only knowledge graph (KG) signals (e.g., KGQE and CL-KGQE) and those integrating KG with document-level graphs (e.g., KAR and QSKG).
4.1. Point of injection
Modern QE methods can be roughly categorized into two levels. On one hand, the query can be implicitly enhanced in the vector space, adjusting its representation to improve retrieval performance. On the other hand, QE can be explicitly performed at the text level, by adding extended words or documents to directly enrich the query content. Clarifying this distinction provides a clear foundation for understanding how and where expansion information can be injected into the retrieval pipeline, which is the focus of this section.As shown in Figure 2, it illustrates the difference in injection points between Implicit QE and Explicit QE.
Figure 2. Application Scenarios of Query Expansion in Information Retrieval
4.1.1. Implicit Embedding-based QE
Setting and notation
Let $q$ be the input query and $\mathcal{R}_{k}=\{d_{1},\dots,d_{k}\}$ the top-$k$ texts returned by a first pass (when available). A document $d$ has a fixed embedding $\mathbf{d}\in\mathbb{R}^{m}$; a query is represented either by a single vector $\mathbf{q}\in\mathbb{R}^{m}$ (bi-encoder) or by a sequence of token embeddings $\{\boldsymbol{\phi}_{q_{i}}\}_{i=1}^{|q|}$ (late interaction). We write inner products as $\langle\mathbf{a},\mathbf{b}\rangle=\mathbf{a}^{\top}\mathbf{b}$ and use $\mathrm{sim}(\cdot,\cdot)$ for cosine similarity when needed. Implicit (embedding-level) QE does not emit new terms; it strengthens the query representation using signals distilled from $\mathcal{R}_{k}$ and then re-queries the unchanged index.
Why implicit QE helps
Compared to generative expansion, implicit QE avoids vocabulary drift and decoding latency. It is especially effective when the first pass already surfaces some on-topic material, because the model can amplify directions in embedding space that correlate with relevance and suppress distracting ones.
Having outlined the basic principles and advantages of implicit QE, we now turn our attention to several representative methods. These approaches illustrate how implicit signals—such as contextual information—can be effectively utilized to enhance query expansion and improve retrieval performance.
ANCE-PRF: refining a single query vector with PRF
ANCE-PRF (Yu et al., 2021) augments a dual-encoder retriever without touching the document index. A dedicated PRF encoder re-encodes the query together with feedback texts:
(12) $\mathbf{q}_{\mathrm{PRF}}=\mathrm{BERT}_{\mathrm{PRF}}\big([\texttt{CLS}]~q~[\texttt{SEP}]~d_{1}~[\texttt{SEP}]~\cdots~[\texttt{SEP}]~d_{k}~[\texttt{SEP}]\big).$
The second-stage score is then the usual inner product $s(q,d)=\langle\mathbf{q}_{\mathrm{PRF}},\mathbf{d}\rangle$, where $\mathbf{d}$ is the fixed document vector from the original ANCE index. Training uses a standard softmax contrastive loss over one positive and sampled negatives. Specifically, let $d^{+}$ represent the positive (relevant) document, and $\mathcal{N}$ denote the set of sampled negative (irrelevant) documents (each $d^{-}\in\mathcal{N}$ is an irrelevant document). The contrastive loss is calculated as:
(13) $\mathcal{L}=-\log\frac{\exp\langle\mathbf{q}_{\mathrm{PRF}},\mathbf{d}^{+}\rangle}{\exp\langle\mathbf{q}_{\mathrm{PRF}},\mathbf{d}^{+}\rangle+\sum_{\mathbf{d}^{-}\in\mathcal{N}}\exp\langle\mathbf{q}_{\mathrm{PRF}},\mathbf{d}^{-}\rangle}.$
Intuition. The encoder learns to attend to feedback tokens that complement the query signal, nudging $\mathbf{q}$ toward regions populated by relevant documents. Cost/risks. One extra encoder forward and one extra nearest-neighbor search at inference; effectiveness depends on $\mathcal{R}_{k}$ containing at least some relevant content.
ColBERT-PRF: injecting discriminative token directions
ColBERT-PRF (Wang et al., 2023b) adapts pseudo-relevance feedback to late interaction (Khattab and Zaharia, 2020). ColBERT scores by matching each query token to its best document token, $s(q,d)=\sum_{i}\max_{j}\langle\boldsymbol{\phi}_{q_{i}},\boldsymbol{\phi}_{d_{j}}\rangle$. ColBERT-PRF enriches the query side with a small set of feedback centroids distilled from $\mathcal{R}_{k}$: (i) collect all token embeddings from $\mathcal{R}_{k}$; (ii) cluster them into $K$ centroids $\{\boldsymbol{\nu}_{\ell}\}_{\ell=1}^{K}$ (where $\boldsymbol{\nu}_{\ell}$ denotes the $\ell$-th centroid vector); (iii) assign each centroid a discriminativeness weight $\sigma_{\ell}=\log\frac{N+1}{N_{\ell}+1}$, where $N_{\ell}$ is the number of passages containing the nearest lexical token to $\boldsymbol{\nu}_{\ell}$ (an IDF proxy) and $N$ is the total number of passages in the collection. The final score appends these centroids as extra “query tokens” with weight $\beta$:
(14) $s_{\mathrm{PRF}}(q,d)=\sum_{i}\max_{j}\langle\boldsymbol{\phi}_{q_{i}},\boldsymbol{\phi}_{d_{j}}\rangle\;+\;\beta\sum_{(\boldsymbol{\nu}_{\ell},\sigma_{\ell})\in F_{e}}\sigma_{\ell}\,\max_{j}\langle\boldsymbol{\nu}_{\ell},\boldsymbol{\phi}_{d_{j}}\rangle,$
where $F_{e}$ contains the top-weighted centroids. Intuition. Feedback centroids capture discriminative semantic directions seen in $\mathcal{R}_{k}$ without committing to specific surface forms, which reduces polysemy drift. Cost/risks. Clustering adds a modest overhead (mitigable with approximate methods). If $\mathcal{R}_{k}$ is off-topic, centroids import noise; selecting small $K$ and using IDF-like weighting helps.
Eclipse: reweighting embedding dimensions with positive and negative signals
Eclipse (D’Erasmo et al., 2025) stays in the single-vector regime and adjusts dimensions rather than tokens. From the first pass it takes a positive set $D^{+}$ (top results) and a negative set $D^{-}$ (tail results). For each embedding dimension $r\in\{1,\dots,m\}$ it estimates an importance
(15) $\mathrm{imp}(r)=\frac{\mathrm{avg}_{d\in D^{+}}\,v_{d,r}-\mathrm{avg}_{d\in D^{-}}\,v_{d,r}}{\sigma_{r}+\epsilon},$
where $v_{d,r}$ is the $r$-th coordinate of document vector $\mathbf{v}_{d}$, $\sigma_{r}$ is the standard deviation of that coordinate across $D^{+}\cup D^{-}$, and $\epsilon$ stabilises the ratio. The augmented query is a Hadamard (elementwise) reweighting, $\mathbf{q}_{\mathrm{aug}}=\mathbf{q}\odot\mathrm{imp}$, and the second pass uses $\langle\mathbf{q}_{\mathrm{aug}},\mathbf{d}\rangle$. Intuition. Dimensions that consistently separate positives from negatives are amplified; dimensions activated by off-topic material are damped. This is a soft, data-driven denoiser in latent space. Cost/risks. No extra encoding, only statistics over $D^{+}$ and $D^{-}$. Choosing $D^{-}$ too aggressively can over-suppress useful but rare signals; small negative pools and standard deviation normalization mitigate this.
QB-PRF: Query-Bag Pseudo-Relevance Feedback (embedding-level)
QB-PRF (Zhang et al., 2024a) augments the query representation using a small set of semantically equivalent query variants selected from an initial retrieval stage, and then fuses them in embedding space. Let $q$ be the input query and $f(\cdot)$ the query encoder used by the retriever. A first pass returns a candidate pool $\mathcal{C}(q)=\{q^{\prime}_{1},\dots,q^{\prime}_{M}\}$ (e.g., related user queries or automatically mined paraphrases).
Selection. Candidates are embedded and filtered for semantic equivalence and lexical diversity to form a query bag $\mathcal{B}(q)\subset\mathcal{C}(q)$. In practice, a contrastive encoder or variational autoencoder scores equivalence; we denote the retained items as $\{q^{\prime}_{i}\}_{i=1}^{n}$.
Fusion. Let $\mathbf{q}=f(q)$ and $\mathbf{q}^{\prime}_{i}=f(q^{\prime}_{i})$. QB-PRF produces an augmented representation by attentional pooling over the bag:
(16) $a_{i}\;=\;\frac{\exp\big(\tau^{-1}\,\mathbf{q}^{\top}\mathbf{q}^{\prime}_{i}\big)}{\sum_{j=1}^{n}\exp\big(\tau^{-1}\,\mathbf{q}^{\top}\mathbf{q}^{\prime}_{j}\big)},\qquad\mathbf{q}_{\text{aug}}\;=\;(1-\beta)\,\mathbf{q}\;+\;\beta\sum_{i=1}^{n}a_{i}\,\mathbf{q}^{\prime}_{i},$
where $\tau>0$ is a temperature and $\beta\in[0,1]$ controls the contribution of the bag. The second pass scores documents with $s(d\mid q^{+})=\mathbf{q}_{\text{aug}}^{\top}\mathbf{d}$ (for dual-encoders) or by replacing the query side in a late-interaction scorer (for multi-representation models).
QB-PRF operates entirely in the latent space: it does not append tokens nor modify the index, and it complements classical PRF by injecting paraphrastic evidence that is robust to lexical mismatch. Reported results show consistent gains in both early precision and deep recall for conversational and ad hoc retrieval settings, with negligible online cost beyond one additional query encoding.
LLM-VPRF: Vector Pseudo-Relevance Feedback for LLM-based retrievers
LLM-VPRF (Li et al., 2025c) extends pseudo-relevance feedback to dense retrievers whose query encoders are derived from large language models, such as PromptReps, RepLLaMA, and LLM2Vec. Given an initial query vector $\mathbf{q}_{0}=f(q)$, the first-pass retrieval returns the top-$k$ documents $\mathcal{R}_{k}(q)=\{d_{1},\dots,d_{k}\}$ with embeddings $\{\mathbf{d}_{j}\}$ and scores $\{s_{j}\}$. The updated query is obtained by combining the original query and the feedback documents:
(17) $\mathbf{q}_{\mathrm{vprf}}=(1-\alpha)\,\mathbf{q}_{0}+\alpha\,\frac{\sum_{j=1}^{k}w_{j}\,\mathbf{d}_{j}}{\sum_{j=1}^{k}w_{j}},$
where $\alpha\in[0,1]$ controls the contribution of feedback, and $w_{j}$ can be uniform, score-based (e.g., softmax over $\{s_{j}\}$), or IDF-based.
The refined query vector is then used for a second retrieval pass on the same index, without text generation, tokenization changes, or model fine-tuning. This approach incorporates corpus-specific semantics from the initial results while remaining lightweight. Experiments on MS MARCO, BEIR, and LoTTE report consistent gains in nDCG@10 and Recall@1000 over zero-shot LLM retrievers, with larger improvements on queries with low first-pass recall, and it can be seamlessly integrated into dual-encoder pipelines alongside fusion or re-ranking methods.
Takeaways
Implicit query expansion offers a lightweight way to strengthen retrieval signals without generating new terms or modifying the index. The main difference across methods lies in where the feedback signal is injected: ANCE-PRF re-encodes the whole query jointly with feedback documents, ColBERT-PRF adds token-like centroids to enrich late interaction, Eclipse rescales embedding dimensions using positive and negative signals, QB-PRF pools over paraphrastic query variants, and LLM-VPRF updates the query vector with document embeddings.
Compared to generative expansion, these methods avoid hallucination and reduce latency, making them suitable when the first-pass retrieval already yields partially relevant results. Their effectiveness ultimately depends on the quality of the initial feedback set, and they are best viewed as precision-preserving, representation-level refinements that can be combined with rerankers or fused with other expansion strategies for robustness.
4.1.2. Selection-based Explicit QE
Problem setup and notation
Let $q$ denote the input query and $\mathcal{R}_{k}=\{D_{1},\dots,D_{k}\}$ the pseudo-relevant set returned by a first pass (e.g., BM25 or a dense retriever). From $\mathcal{R}_{k}$ we derive a pool of candidate units $C$ to consider for expansion. A candidate can be a term $t$, an $n$-gram, or a short chunk $c$. We write $\text{sim}(\mathbf{a},\mathbf{b})$ for cosine similarity between vectors $\mathbf{a}$ and $\mathbf{b}$, and $\langle\mathbf{a},\mathbf{b}\rangle$ for an inner product. Selection-based explicit QE uses a pretrained language model (PLM) not to generate text, but to score candidates in context and keep only those that are likely to help retrieval; the chosen items are then inserted into the query or used as additional evidence at scoring time. This design gives tight control over vocabulary and makes it easy to incorporate domain or user constraints.
The core idea of selection-based explicit QE with PLMs is to identify salient terms or chunks from the input query and incorporate them into the reformulated query. Building on this principle, several representative approaches have been proposed, which differ in how they select informative units and integrate them with the original query. In the following, we present some of these representative methods in detail.
CEQE: Context-aware scoring in an RM3 pipeline
CEQE (Naseri et al., 2021) replaces frequency counts in RM3 (a relevance model variant that performs linear interpolation of expansion terms with the original query using the Query Likelihood score) with context-sensitive similarities computed by BERT. First, retrieve $\mathcal{R}_{k}$ and encode the query to a vector $\mathbf{q}$. For each document $D\in\mathcal{R}_{k}$ and each occurrence of a term $t$ in $D$, obtain the contextual embedding $\mathbf{m}^{D}_{t}$ from an intermediate BERT layer. Let $\mathbf{m}^{D}_{*}$ denote the set of all contextual embeddings of all terms in $D$ The contribution of $t$ in $D$ is the normalised similarity
(18) $p(t\mid q,D)\;=\;\frac{\sum_{\mathbf{m}^{D}_{t}}\text{sim}(\mathbf{q},\mathbf{m}^{D}_{t})}{\sum_{\mathbf{m}^{D}_{*}}\text{sim}(\mathbf{q},\mathbf{m}^{D}_{*})}.$
Aggregate over feedback documents with an RM3-style weighting $p(q\mid D)$ (the Query Likelihood score of $q$ given $D$) to obtain a feedback model
(19) $p(t\mid\theta_{\mathcal{R}})\;\propto\;\sum_{D\in\mathcal{R}_{k}}p(t\mid q,D)\,p(q\mid D).$
Finally, interpolate the top-$m$ terms with the original query using a weight $\lambda$. Intuitively, a term that appears near query-relevant contexts (even if rare) is promoted, while spurious high-frequency terms are down-weighted. CEQE is fully unsupervised and drops into existing PRF pipelines with only additional BERT text encodes.
SQET: Supervised filtering of term candidates
SQET (Naseri et al., 2022) turns term selection into a learned binary decision. Given $q$ and a candidate $t$ (with or without a short context window), a cross-encoder processes the pair and outputs a relevance probability $s(t\mid q)\in[0,1]$. When multiple mentions are available, instance-level scores $\{s_{i}\}$ are combined into a term-level score using simple rules:
(20) $\text{Max: }\max_{i}s_{i},\qquad\text{Weighted-sum: }\frac{1}{Z}\sum_{i}\text{tf}(t;\text{ctx}_{i})\,s_{i},\qquad\text{invRank: }\frac{1}{Z}\sum_{i}\frac{s_{i}}{\log_{2}(\text{rank}(\text{ctx}_{i})+1)}.$
The top-ranked terms are interpolated with the original query exactly as in PRF. Because the model observes query–term pairs during training, SQET better rejects topical but off-target terms than purely unsupervised scoring.
BERT-QE: Selecting chunks as expansion evidence for re-ranking
BERT-QE (Zheng et al., 2020) selects short text chunks from $\mathcal{R}_{k}$ and reuses them as evidence when re-ranking candidates. Pipeline: (1) obtain $\mathcal{R}_{k}$ and, if desired, re-rank it once with a cross-encoder to improve the feedback pool; (2) split each $D\in\mathcal{R}_{k}$ into overlapping chunks $\{c_{i}\}$ and score $\text{rel}(q,c_{i})$ with the cross-encoder; (3) keep the top-$k_{c}$ chunks $C$ and, for each candidate document $d$, compute an evidence score
(21) $\text{rel}(C,d)\;=\;\sum_{c_{i}\in C}\pi_{i}\,\text{rel}(c_{i},d),\qquad\pi_{i}=\frac{\exp(\text{rel}(q,c_{i}))}{\sum_{c_{j}\in C}\exp(\text{rel}(q,c_{j}))}.$
The final score interpolates the original and evidence components:
(22) $\text{rel}(q,C,d)=(1-\alpha)\,\text{rel}(q,d)+\alpha\,\text{rel}(C,d).$
Rather than adding new tokens to $q$, BERT-QE carries forward small, query-vetted excerpts that the re-ranker can match against candidate documents, yielding robust gains with a controllable compute budget.
CQED: Domain-aware term selection for biomedical search
CQED (Khader and Ensan, 2023) targets scholarly biomedical search by combining general BERT with UMLS-BERT and lightweight preprocessing. Named entities in $q$ are first identified and protected to avoid drift. A masking module creates masked variants of $q$ to elicit semantically close substitutes. Each mask is filled by both encoders to form a candidate pool $C=\{t\}$ that mixes general and domain-specific vocabulary. A pairwise learning-to-rank model then scores $t\in C$ by the improvement it brings when appended to $q$; multi-head attention fuses signals from the two encoders before a final scoring layer. The expanded query $q^{+}=q\cup\text{Top-}k(C)$ is used for retrieval. By anchoring term proposals in biomedical knowledge while keeping a general encoder in the loop, CQED reduces off-domain expansions and has reported strong gains on TREC-COVID.
PQEWC: Personalizing term choices from a user’s history
PQEWC (Bassani et al., 2023) personalizes explicit expansion using the user’s past interactions. From clicked titles or prior reads for user $u$, the method builds an embedding pool $E_{u}=\{\mathbf{w}^{(u)}\}$ with a ColBERT encoder. Two selection strategies are considered. Cluster-based: cluster $E_{u}$ (e.g., HDBSCAN), rank cluster centroids by similarity to the current query vector, and pick the nearest token to each top centroid as expansion terms. Fast approximate: directly retrieve the $n$ nearest tokens to the query vector from $E_{u}$ using an ANN index. Each selected term $t_{i}$ receives a weight $w_{i}$ proportional to its similarity to $q$, and the scoring interpolates original and personalized terms:
(23) $\text{Score}(q^{\prime},d)=(1-\alpha)\,\text{Score}(q,d)+\alpha\sum_{i}w_{i}\,\text{Score}(t_{i},d).$
This design steers ambiguous or broad queries toward a user’s long-term interests at millisecond-level overhead.
Design notes
Selection-based explicit QE keeps the expansion vocabulary under control, is easy to constrain with domain ontologies or user profiles, and plays well with both sparse and neural rankers. Unsupervised scoring (CEQE) is simple and strong; supervised filtering (SQET) improves precision when labels exist; chunk selection (BERT-QE) lets a re-ranker exploit richer contextual evidence; domain and personalization layers (CQED, PQEWC) target drift and ambiguity in practice.
4.2. Grounding and interaction
Global notation
Throughout this subsection, unless otherwise noted: $q$ denotes the original input query, and $g$ a generated expansion (e.g., a passage, rationale, pseudo-document, or clue). $G(q)=\{g_{1},\dots,g_{m}\}$ is the set of generated expansions. Concatenation at the string level is $[q;g]$, and an augmented query is denoted $q^{+}$ (for either textual or vector-space form). $R_{k}(q)=\{d_{1},\dots,d_{k}\}$ represents the top-$k$ pseudo-relevant documents from a retriever, scored by $s(q,d)$. $\mathbf{q}$ and $\mathbf{d}$ are vector embeddings of $q$ and $d$, and $\mathrm{rank}(d\mid R(\cdot))$ gives the position of $d$ in the ranked list.
4.2.1. Zero-Grounding, Non-Interactive QE: LLM-Driven Single-Stage Expansion
Scope and specific notation
This category comprises zero- or few-shot query expansion methods that rely solely on prompting a large language model ($\mathcal{L}$) without supervised fine-tuning, additional retriever training, or systematic grounding in the target corpus. Minor uses of corpus evidence in control variants (e.g., injecting a few top-ranked passages into the prompt) are still considered in scope when the main expansion process remains prompt-only and non-interactive. Beyond the global notation, we write $\overline{\mathbf{v}}$ for the mean embedding over multiple $g\in G(q)$, used in vector aggregation, and $p(w\mid\theta_{\bullet})$ for a relevance model estimated from generated feedback.
Query2Doc: Few-shot pseudo-documents
Q2D (Wang et al., 2023c) prompts $\mathcal{L}$ to write a short passage that would plausibly answer $q$ using $k$ in-context exemplars (e.g., k=4k{=}4). The pseudo-document $g$ is then concatenated with $q$ to obtain $q^{+}=[q;g]$. For sparse retrieval, repeating the original query tokens before concatenation (e.g., five repetitions) preserves term salience; for dense models, $q$ and $g$ are joined with a separator and encoded jointly or separately. Q2D reliably improves BM25 on MS MARCO and TREC DL (e.g., up to $15.2\%$ nDCG@10 on DL’20) and gives modest gains when used to augment dense training. Trade-offs include single-shot hallucinations and non-trivial decoding latency.
CoT-QE: Reason-then-expand prompting
CoT-QE (Jagerman et al., 2023) elicits an explanatory chain before the answer with a zero-shot prompt of the form: “Answer the following query: {q}. Give the rationale before answering.” The produced rationale $g$ contains definitions, disambiguation cues, and facet hints; $[q;g]$ is issued to the retriever. A PRF-enhanced variant (CoT/PRF) injects the top-3 BM25 passages into the prompt to bias reasoning toward corpus-specific terminology. CoT-QE outperforms classical PRF baselines (e.g., Bo1, KL) and other prompt baselines on MS MARCO/BEIR, especially on top-heavy metrics (MRR@10, nDCG@10), and remains effective with smaller models.
GAR: Generation-Augmented Retrieval
GAR (Mao et al., 2020) trains a seq2seq model (BART-large) on open-domain QA pairs to emit one of three query-specific strings: an answer, a declarative sentence containing the answer, or the title of a relevant passage.
At inference, the generated string $g$ is concatenated to the original query, yielding $[q;g]$, which is sent to BM25 or a dense retriever. Running the three generators produces three rankings that can be combined by simple rank fusion for additional recall. GAR is attractive because it is plug-and-play and task-aligned (the model learns to verbalise what a good answer looks like), but it pays generation latency and can drift if $g$ over-commits to a specific interpretation.
GRF: Generative Relevance Feedback without First-Pass Dependence
GRF (Mackie et al., 2023b) replaces $R_{k}(q)$ with an LLM-produced set $G(q)$ and estimates a relevance model directly over $G(q)$. Concretely, the RM3-style term distribution is
(24) $p(w\mid\theta_{G})\propto\sum_{g\in G(q)}p(w\mid g)\,p(q\mid g),$
where $p(w\mid g)$ comes from term statistics of $g$ and $p(q\mid g)$ is approximated by a language-model likelihood. GRF prompts for corpus-matched genres (keywords, entities, pseudo-documents, news-style summaries), which aligns expansion with collection discourse and yields sizable MAP and nDCG@10 gains on Robust04, CODEC, and TREC DL.
GRF for dense and learned-sparse, and fusion with PRF. A follow-up study applies GRF to dense and learned-sparse retrievers by encoding each $g\in G(q)$ and aggregating with the original query embedding via a Rocchio-style update; for learned sparse, term weights from $G(q)$ are pruned before indexing (Mackie et al., 2023a). Query-level analyses show complementarity with classical PRF: GRF helps low-recall queries; PRF strengthens already-recalled facets. A weighted reciprocal-rank fusion of GRF and PRF improves Recall@1000 in the majority of cases.
HyDE: Hypothetical document embeddings
HyDE (Gao et al., 2023) prompts $\mathcal{L}$ to draft one or more short, plausible passages answering $q$ (e.g., “write a paragraph that answers the question”), then uses their embeddings as surrogates for the missing context in zero-shot dense retrieval. Let $G=\{g_{1},\dots,g_{N}\}$ be generated passages and $f(\cdot)$ the fixed dense encoder. Form a pseudo-context vector $\overline{\mathbf{v}}=\frac{1}{N}\sum_{i=1}^{N}f(g_{i})$ and update the query embedding by a convex combination
(25) $\tilde{\mathbf{q}}=(1-\alpha)\,f(q)+\alpha\,\overline{\mathbf{v}},\qquad\alpha\in[0,1].$
Retrieving with $\tilde{\mathbf{q}}$ improves zero-shot dense baselines on heterogeneous benchmarks by bridging the representation gap when no relevance labels or in-domain training are available; the main risks are drift from fanciful generations and added encoding cost for multiple $g_{i}$.
Exp4Fuse: Rank fusion for LLM-augmented sparse retrieval
Exp4Fuse (Liu and Zhang, 2025) prompts an LLM in a zero-shot setting to generate a single hypothetical document $r_{q}$. The original query $q_{o}$ is repeated $\lambda$ times for balance and concatenated with $r_{q}$:
(26) $q_{e}=\big[\,\underbrace{q_{o},\dots,q_{o}}_{\lambda\ \mathrm{times}},r_{q}\,\big],\quad\lambda=5.$
Using the same sparse retriever (e.g., SPLADE++), two ranked lists are obtained: $I_{oq}$ from $q_{o}$ and $I_{eq}$ from $q_{e}$. A modified reciprocal-rank fusion then scores each document $d$ as:
(27) $s_{\mathrm{FR}}(d)=\left(w_{i}+\frac{n}{10}\right)\sum_{i=1}^{2}\frac{1}{k+r_{i}},$
where $k=60$ mitigates outliers, $n\in\{1,2\}$ counts the number of list appearances, $w_{i}=1$ weights routes equally, and $r_{i}$ is the rank of $d$ in list $i$ (or $\infty$ if absent). This indirect QE boosts nDCG@10 and Recall@1000 for learned sparse models on MS MARCO and BEIR, often achieving SOTA with low overhead; regressions on strong queries suggest selective invocation via query-quality predictors.
Contextual clue sampling and fusion
Contextual clue sampling (Liu et al., 2022) targets breadth and control. A generator produces many short, answer-oriented “clues” for $q$ using stochastic decoding. Candidates are clustered (e.g., by edit distance), keeping one representative per cluster to reduce redundancy. Each retained clue $c_{i}$ forms an expanded query $q;c_{i}]$ and yields a ranked list with score $s([q;c_{i}],d)$. Final ranking fuses these lists with generation-aware weights:
(28) $s_{\mathrm{fused}}(q,d)=\sum_{i=1}^{M}P_{\mathrm{gen}}(c_{i}\mid q)\;s([q;c_{i}],d).$
This scheme encourages multiple reasoning paths while placing more trust in clues the model deems probable. It is particularly effective for multi-hop or under-specified questions.
SEAL: Generating substrings as executable keys
SEAL (Bevilacqua et al., 2022) constrains generation to $n$-gram substrings that are guaranteed to occur in the target corpus. A BART model is fine-tuned to produce such substrings, and decoding is enforced with an FM-index (a compressed full-text substring index) so that every token sequence $t$ is corpus-valid. Each candidate substring $t$ is scored by a rarity-adjusted language-model score that balances ”query plausibility” (from the BART model) and ”corpus rarity” (from the FM-Index):
(29) $s_{\mathrm{SEAL}}(t\mid q)=\log P_{\mathrm{LM}}(t\mid q)\;-\;\lambda\,\log P_{\mathrm{corpus}}(t),$
where $P_{\mathrm{corpus}}(t)$ is the prior probability of $t$ in the target corpus, calculated as $\frac{F(t,\mathcal{R})}{\sum_{d\in\mathcal{R}}|d|}$—here, $\mathcal{R}$ represents the target retrieval corpus, and $F(t,\mathcal{R})$ is the total number of occurrences of $t$ in $\mathcal{R}$; $\lambda$ is an interpolation hyperparameter that balances the two terms. This scoring formula biases toward informative (rarer) substrings that remain plausible under $q$. Documents containing high-scoring substrings are retrieved via FM-index lookups, and evidence from multiple substrings is aggregated. SEAL unifies expansion and retrieval: generated outputs are directly executable keys, avoiding hallucination and enabling GPU-free, low-latency search with strong accuracy on KILT-style benchmarks.
This family covers methods that condition expansion on retrieval evidence and/or run multiple retrieve–expand–requery cycles. Compared with free-running prompt-only generation, these approaches explicitly close the loop with the collection, which improves grounding, controllability, and robustness.
4.2.2. Grounding-Only, Non-Interactive QE: Corpus-Evidence Anchored Single-Pass Expansion
Scope and specific notation
This category covers retrieval-conditioned expansion methods in which $\mathcal{L}$ is explicitly grounded in the target collection during a single retrieve–generate–requery pass. Grounding can involve: (i) conditioning on pseudo-relevant passages $R_{k}(q)$; (ii) constraining generation to corpus-valid substrings or entities; or (iii) applying selection/calibration based on first-pass retrieval scores. Compared with Zero-Grounding QE, these methods access corpus evidence; unlike interactive QE, they do not loop over multiple passes. Beyond the global notation, we introduce $P_{\mathrm{gen}}(g\mid q,R_{k}(q))$ as the generation probability given corpus context, and $\lambda\in[0,1]$ as the interpolation parameter for fusion between retrieval runs.
Practical notes
Retrieval-conditioned QE methods vary in how they incorporate collection evidence. Constraint-based approaches such as SEAL ensure grounding by limiting outputs to corpus-valid substrings, eliminating hallucination and reducing latency. Selection-based methods like EAR filter diverse generations through learned rerankers, while feedback-calibrated methods such as MUGI and FGQE adjust expansion impact using retrieval signals or fairness objectives. Hybrid schemes, e.g., GenPRF and PromptPRF, combine statistical PRF signals with LLM-generated semantics to balance recall and precision. CSQE mixes grounded evidence from $R_{k}(q)$ with unguided world-knowledge generations to increase robustness. Overall, effective designs couple generative expressiveness with reliable grounding—via hard constraints, learned selection, feedback calibration, or explicit interpolation with traditional signals.
MILL: Mutual verification with corpus evidence
MILL  (Jia et al., 2023) constructs two complementary contexts and keeps only what they agree on. Stage 1 generates multiple explanations by a query–query–document style prompt (decompose $q$ into sub-queries and write short passages), and in parallel retrieves pseudo-relevant documents with BM25. Stage 2 performs mutual verification: generated texts are scored by their semantic consistency with retrieved documents (cosine between encoder representations), and PRF documents are rescored by their agreement with the generated explanations. Only top-agreement items from both sides are retained, concatenated to $q$, and reissued. Across TREC-DL’19/20 and BEIR, MILL improves NDCG@1000, MRR@1000, and Recall@1000, with largest gains in specialized domains, indicating that agreement filtering curbs hallucination while preserving useful diversity.
AGR: Analyze–Generate–Refine
AGR (Chen et al., 2024) structures prompting into three passes. Analyze: extract key phrases and a brief need statement for $q$. Generate: produce several answer-oriented candidates and, for each, run a light retrieval to fetch a small set of supporting passages; ask $\mathcal{L}$ to enrich each candidate using that context. Refine: perform a self-review over all enriched candidates to remove off-topic or redundant content and return a compact expansion $g$. The final $[q;g]$ improves zero-shot retrieval and end-to-end OpenQA EM across NQ, TriviaQA, WebQ, and CuratedTREC, showing that staged quality control reduces error propagation without supervision.
EAR: Expand, rerank, and retrieve
EAR (Chuang et al., 2023) decouples generation from selection. A generator (e.g., BART-large or T0-3B) first proposes a diverse set of candidate expansions $E=\{e_{1},\dots,e_{N}\}$. A learned reranker then predicts the utility of each candidate. Two variants are common: a retrieval-independent reranker that scores pairs $(q,e)$, and a retrieval-dependent reranker that also conditions on a top passage $(p_{1})$, scoring triples $(q,e,p_{1})$. With a pairwise loss, the reranker learns to prefer expansions that raise the rank of the gold passage under BM25. The best candidate $\hat{e}$ is concatenated with the query and used for the final retrieval. By filtering the generator through a PLM-based selector, EAR preserves diversity while curbing noisy $e_{i}$, delivering consistent gains over single-shot generators.
GenPRF: Generative pseudo-relevance feedback
GenPRF (Wang et al., 2023a) marries classical RM3 with neural reformulation. From the top-ranked texts of a first pass, short passages are selected (e.g., FirstP/TopP/MaxP) and fed to a seq2seq model (T5 or FLAN-T5) to produce a corpus-aware reformulation $q_{r}$. Rather than replacing RM3, GenPRF mixes the two query models. Let $p(w\mid\theta_{\mathrm{RM3}})$ be RM3’s term distribution and $p(w\mid\theta_{\mathrm{Gen}})$ the term distribution estimated from $q_{r}$ (e.g., normalised token weights). GenPRF interpolates
(30) $p(w\mid\theta^{\dagger})=k_{\mathrm{RM3}}\,p(w\mid\theta_{\mathrm{RM3}})+k_{\mathrm{Gen}}\,p(w\mid\theta_{\mathrm{Gen}}),\qquad k_{\mathrm{RM3}}+k_{\mathrm{Gen}}=1,$
and uses $p(\cdot\mid\theta^{\dagger})$ in the second pass. This hybrid retains RM3’s high-recall term statistics while injecting semantically richer paraphrases from the generator, yielding consistent MAP and nDCG gains across ad hoc collections.
CSQE:Corpus-Steered query expansion
CSQE (Lei et al., 2024)balances LLM world knowledge with collection specificity by combining two sources: (i) sentences extracted from $R_{k}(q)$ that an LLM deems query-relevant and (ii) a hypothetical document generated without corpus input. Concatenating both with $q$ improves robustness on TREC DL and BEIR and excels on NovelEval, where the target knowledge is absent from LLM pretraining.
MUGI:Multi-text generation integration for query expansion
MUGI (Zhang et al., 2024b) prompts an LLM in a zero-shot setting to generate multiple pseudo-reference documents $R=\{r_{1},\dots,r_{n}\}$ (e.g., n=5n{=}5), which are then integrated with the original query $q$ to enrich both sparse and dense retrieval without additional training. For sparse retrieval (BM25), the pseudo-references are concatenated with repeated copies of $q$ to balance their influence, given BM25’s sensitivity to term frequency and document length. The repetition factor $\lambda$ is adapted to reference length:
(31) $q_{\mathrm{sparse}}=\big[\,\underbrace{q,\dots,q}_{\lambda\ \text{times}},\ r_{1},\dots,r_{n}\,\big],$
where $\lambda$ is a length-normalized integer (hyperparameters described in the appendix). For dense retrieval, pairwise concatenations $\big[\,q,r_{i}\,\big]$ are encoded independently and pooled:
(32) $e_{q}=\frac{1}{n}\sum_{i=1}^{n}f\big([\,q,r_{i}\,]\big),$
yielding a query embedding enriched with multiple contextualized references. A light Rocchio-style calibration shifts the query embedding toward high-agreement positives (generated references and top-$K$ documents) and away from low-ranked negatives, mitigating noisy generations. This unified approach improves nDCG@10 and Recall@1000 across sparse and dense models on TREC DL and BEIR, enabling smaller LLMs to match or surpass larger models.
PromptPRF:a feature-based PRF framework
Rather than free-form text, PromptPRF (Li et al., 2025b)extracts structured features from $R_{k}(q)$ using an LLM—keywords, entities, focused summaries, and chain-of-thought-derived terms—and encodes the result once as an augmented query representation for the second pass. Because features are grounded in PRDs and can be produced offline, PromptPRF reduces online LLM cost while outperforming term-statistics PRF baselines.
FGQE:Fair Generative Query Expansion
FGQE (Jaenich et al., 2025)measures exposure disparities in the first-pass ranking using Average Weighted Rank Fairness and then conditions the LLM to generate entities, keywords, or pseudo-documents that target underexposed groups. The resulting expansions improve fairness with minimal loss in nDCG, illustrating that feedback signals can optimize objectives beyond relevance.
4.2.3. Grounding-Aware Interactive QE: Multi-Round Retrieve-Expand Loops
Scope and specific notation
This category covers iterative query expansion methods that interleave retrieval and generation over $T>1$ rounds, with at least some expansions grounded in documents retrieved in earlier stages. This closed-loop setting enables progressive query refinement, going beyond the single-pass grounding of the previous category. Building on the global notation, let the initial query be $q^{(1)}=q$. At iteration $t$ ($1\leq t\leq T$), the model produces an expansion $g^{(t)}$ (from prior knowledge or conditioned on $R_{k}(q^{(t)})$) and forms:
(33) $q^{(t)}_{\mathrm{exp}}=[\,q^{(t)};g^{(t)}\,].$
Retrieved sets $R_{k}(q^{(t)})$ may be combined, filtered, or re-used in later iterations, and stopping criteria may depend on retrieval budgets or observed gains.
InteR:a novel framework that facilitates information refinement through synergy between RMs and LLMs
InteR (Feng et al., 2023) alternates expansion and retrieval. An LLM first produces expansion text to pull in an initial set; those passages are then summarized into a knowledge collection $S$ that is interleaved with $q$ to requery. The method supports hybrid sparse–dense stacks and shows strong results on large-scale and low-resource search.
ProQE:progressive query expansion
When retrieval calls are expensive, ProQE (Rashid et al., 2024) adopts a progressive loop that retrieves only one document per iteration, lets an LLM judge and extract terms, updates the query, and repeats until a budget is met. A final chain-of-thought expansion is added before a single high-recall sweep, delivering higher effectiveness than single-shot PRF or zero-shot LLM QE under tight budgets.
LameR:the Language language model as Retriever
LameR  (Shen et al., 2023)performs a small first-pass retrieval, then prompts the LLM to produce multiple likely answers conditioned on those few passages, and finally issues an answer-augmented query for the second pass. This answer-first strategy improves recall without retriever fine-tuning and sets competitive zero-shot baselines on DL’19/DL’20 and BEIR.
ThinkQE: Iterative thinking with corpus interaction
ThinkQE (Lei et al., 2025) alternates explicit reasoning and retrieval. At round $t$, the model lists interpretations and related concepts, then writes a short expansion $g_{t}$; the expanded query $q;g_{1};\dots;g_{t}]$ is issued, and only new top-$k$ documents are fed back as context for the next round to encourage coverage. After a small number of rounds, all expansions are concatenated and searched. On TREC DL and BRIGHT, this retrieve–expand–filter loop increases both nDCG@10 and Recall@1000 versus single-shot prompting, suggesting that iterative, diversity-seeking reasoning is an effective antidote to early bias.
4.2.4. Takeaways
Across the grounding–interaction spectrum, LLM-based query expansion trades between prior-knowledge breadth and corpus-specific precision. Zero-grounding methods (e.g., Q2D, CoT-QE, HyDE) are simple to deploy and can lift recall without first-pass cost, but risk drift and hallucination without corrective signals. Grounding-only, non-interactive designs (e.g., SEAL, MILL, EAR, GenPRF, PromptPRF) add a light first-pass to curb hallucination, align terminology, and enable richer control via constraints, rerankers, or interpolation with statistical PRF—often delivering stable gains for both sparse and dense retrievers. Grounding-aware interactive loops (e.g., InteR, ProQE, ThinkQE) further close the feedback loop, progressively refining expansions to recover missed facets, at higher latency and engineering cost. Effective practice couples generative expressiveness with targeted grounding and, where feasible, iterative refinement; parameter choices ($k$, $\lambda$, iteration budget) and gating strategies determine the cost–quality trade-off in production.
4.3. Learning and Alignment for QE with PLMs/LLMs
Scope and notation
This section covers methods that explicitly align large language models or retrievers to produce query expansions that improve downstream retrieval. We use $q$ for a query, $g$ for an expansion text, $d^{+}$ for a gold (relevant) passage, and $D^{-}$ for a set of negatives. Encoders map text to vectors: $f(\cdot)$ denotes the student retriever, $f_{T}(\cdot)$ a teacher retriever. When needed, $\mathcal{L}$ denotes an LLM used to generate $g$. For dense training, we write a standard contrastive (InfoNCE) objective
(34) $\mathcal{L}_{\mathrm{cont}}(q)=-\log\frac{\exp(\,\mathrm{sim}(f(q),f(d^{+}))/\tau\,)}{\exp(\,\mathrm{sim}(f(q),f(d^{+}))/\tau\,)+\sum_{d^{-}\in D^{-}}\exp(\,\mathrm{sim}(f(q),f(d^{-}))/\tau\,)},$
with temperature $\tau$ and cosine similarity $\mathrm{sim}(\cdot,\cdot)$. Supervised fine-tuning (SFT) minimizes negative log-likelihood of a target expansion sequence $g$ under a generator $\pi_{\theta}$: $\mathcal{L}_{\mathrm{SFT}}=-\log\pi_{\theta}(g\mid q)$. Direct Preference Optimization (DPO) aligns a generator $\pi_{\theta}$ to prefer $g^{+}$ over $g^{-}$ for the same $q$:
(35) $\mathcal{L}_{\mathrm{DPO}}=-\log\sigma\!\Big(\beta\big[\log\pi_{\theta}(g^{+}\!\mid\!q)-\log\pi_{\theta}(g^{-}\!\mid\!q)-\log\pi_{\mathrm{ref}}(g^{+}\!\mid\!q)+\log\pi_{\mathrm{ref}}(g^{-}\!\mid\!q)\big]\Big),$
with inverse temperature $\beta$ and a frozen reference model $\pi_{\mathrm{ref}}$. Parameter-efficient fine-tuning (PEFT) such as LoRA can be used in any generator or encoder below to reduce memory and time.
4.3.1. SoftQE: distilling LLM-generated expansions into query representations
SoftQE (Pimpalkhute et al., 2024) converts the benefit of LLM expansions into a cheaper retriever that no longer requires the LLM at inference. Offline, a large language model (e.g., GPT-3.5-turbo) generates a pseudo-document $g$ for each $q$ (e.g., Query2Doc-style). A teacher dual-encoder (denoted as $f_{T}(\cdot)$) is trained on expanded inputs $[q;g]$ with contrastive loss, producing a strong representation $f_{T}([q;g])$. A student encoder (denoted as $f(\cdot)$) then learns, from only the raw query $q$, to approximate the teacher’s expanded representation via an MSE distillation term,
(36) $\mathcal{L}_{\mathrm{dist}}=\big\|f(q)-f_{T}([q;g])\big\|_{2}^{2},$
combined with the usual contrastive objective on raw queries:
(37) $\mathcal{L}_{\mathrm{SoftQE}}=\alpha\,\mathcal{L}_{\mathrm{dist}}+(1-\alpha)\,\mathcal{L}_{\mathrm{cont}}(q).$
Intuitively, the student is taught to embed $q$ as if the helpful expansion were already present. The result is an expansion-aware retriever with the latency of a standard encoder. In practice, SoftQE yields sizable nDCG@10 gains on MS MARCO and BEIR while avoiding on-the-fly decoding costs; adding cross-encoder distillation further sharpens decision boundaries.
4.3.2. RADCoT: retrieval-augmented distillation of chain-of-thought expansions
RADCoT (Lee et al., 2024) compresses a reasoning-capable teacher (e.g., GPT-3) into a smaller retrieval-augmented student that can produce chain-of-thought (CoT) style expansions efficiently. Stage one collects pairs $(q,g^{\mathrm{CoT}})$ by prompting the teacher to explain and answer $q$. Stage two trains a Fusion-in-Decoder (FiD) student conditioned on retrieved contexts $P=\{p_{1},\dots,p_{K}\}$, minimizing the likelihood loss
(38) $\mathcal{L}_{\mathrm{RADCoT}}=-\log P_{\mathrm{FiD}}\!\big(g^{\mathrm{CoT}}\mid q,P\big).$
At inference, the student generates a short rationale $g^{\mathrm{CoT}}_{\mathrm{stu}}$ grounded in the passages and appends it to $q$. Because the student relies on retrieval rather than memorized world knowledge, it achieves teacher-like expansion quality with orders-of-magnitude fewer parameters; using PEFT (e.g., LoRA adapters) is sufficient for stable training.
4.3.3. ExpandR: alternating retriever training and preference-aligned generation
ExpandR (Yao et al., 2025) jointly optimizes a retriever and an LLM so that expansions are both informative and useful to the ranker. The procedure alternates two steps.
Retriever step. Given $q$ and an LLM expansion $g$, construct a query representation by averaging the embeddings of the original query and the expanded query (gating is optional for weight adaptation):
(39) $\tilde{q}=\frac{1}{2}\left(f(q)+f([q;g])\right)$
where $f(\cdot)$ denotes the query encoder, and $\tilde{q}$ is the augmented query embedding. The retriever is then trained with $\mathcal{L}_{cont}(q)$ to learn to exploit the additional semantic signal from $g$.
Generator step. Freeze the retriever and update the LLM with DPO using pairs $(g^{+},g^{-})$ scored by a reward that combines (i) self-consistency with the LLM’s own answer conditioned on $d^{+}$ and (ii) retrieval utility, e.g., the rank of $d^{+}$ under $[q;g]$. The DPO loss nudges the LLM toward expansions that lift the actual ranking objective without requiring a separate learned reward model.
By repeating these steps, the retriever becomes expansion-aware while the generator becomes retriever-aware, producing generalizable gains across MS MARCO and BEIR, especially on collections with large lexical–semantic gaps.
4.3.4. Aligned Query Expansion: SFT and DPO for retrieval-oriented generation
AQE (Yang et al., 2025) starts from a base LLM and explicitly reshapes its output distribution toward retrieval-helpful expansions. For each $q$, the base model proposes a pool of candidates $\{g_{i}\}_{i=1}^{n}$; each candidate is scored by how highly $d^{+}$ ranks when searching with $q;g_{i}]$. AQE supports two alignment regimes.
RSFT. Keep only top candidates as targets and run supervised fine-tuning with the token-level loss $\mathcal{L}_{\mathrm{SFT}}$. This anchors the model on high-utility expansions.
DPO. Form preference pairs $(g^{+},g^{-})$ from the top and bottom of the pool and optimize $\mathcal{L}_{\mathrm{DPO}}$ against a frozen reference. DPO is a light-weight alternative to RLHF that avoids training a separate reward model while still encoding utility-aware preferences.
At inference, AQE generates a single short $g$ via greedy decoding and concatenates it to $q$, eliminating generate–rank–filter loops. On MS MARCO and NQ, both RSFT and DPO variants improve nDCG@10 and Recall@1000; combining RSFT (to stabilize behavior) followed by DPO (to sharpen preferences) is often strongest. PEFT substantially reduces fine-tuning cost without hurting effectiveness.
Implementation notes and takeaways
(i) Alignment converts brittle prompt-only QE into a reliable component: the generator learns what helps the ranker, and the retriever internalizes expansion effects to amortize decoding cost. (ii) Distillation paths (SoftQE, RADCoT) are preferable when inference budget is tight or privacy rules preclude online LLM calls. (iii) Preference optimization (AQE, ExpandR) directly ties generation quality to ranking gains and is robust across domains; DPO is a pragmatic substitute for full RLHF. (iv) In all regimes, PEFT (LoRA/adapters) makes alignment feasible on modest hardware, and caching expansions or training expansion-aware students keeps serving costs close to vanilla retrieval.
4.4. KG-augmented Query Expansion
Scope and specific notation
[…]
It should be noted that KG-augmented QE here is not independent of the ”grounding” framework in Sec.4.2: Sec.4.2 defines ”grounding” as ”expansion signals relying on real-time evidence from the current retrieval corpus”, while this section takes external KGs as a new grounding source, forming a complementation of ”corpus grounding (Sec.4.2) + KG grounding (this section)”. Corpus grounding ensures alignment with the target corpus, while KG grounding supplements structured semantics missing in the corpus, jointly addressing complex retrieval needs.
After introducing the basic concepts of knowledge graphs, we note that the key factor in KG-augmented query expansion lies in how the extracted entities are utilized. Existing approaches can be broadly divided into two categories: (i) leverages knowledge graph retrieval to incorporate related entities and relations.(ii) links the extracted entities to external corpora or text chunks to obtain additional expansion information.
4.4.1. Entity-based Expansion via Knowledge Graph Retrieval
Scope and notation
In this category of methods, entities are first extracted from the query, and then their neighboring nodes or relations in the knowledge graph are retrieved to enrich the query representation. The expanded query can then be expressed as:
(40) $q^{+}=\big[q\;;\;\bigcup_{e\in\mathcal{E}_{q}}f(\mathcal{N}(v),\mathrm{R}(v))\big],$
where $f(\cdot)$ is a function that selects or weights expansion terms from neighboring entities and relations, and $q;\cdot]$ denotes concatenation or fusion with the original query. In the following, we present specific approaches for Entity-based Expansion.
Entity-aware text injection (KGQE)
KGQE (Perna, 2025) injects compact KG facts directly into the query text to improve disambiguation and coverage. First, entities are detected and linked, $\mathcal{E}_{q}=\mathrm{EL}(q)$. Then short textual snippets for each $e\in\mathcal{E}_{q}$ (e.g., label, type, canonical alias) are selected and concatenated:
(41) $q^{+}\;=\;\mathrm{concat}\Big(q;\;\big[\,\text{text}(e)\,\big]_{e\in\mathcal{E}_{q}}\Big).$
Optional guards include type filters (retain only domain-relevant types), length caps, and de-duplication. The enriched query is issued to the retriever, or used to prompt an LLM when a downstream generator is present. A minimal prompt template is:
  QUERY: "When is the last episode of season 8 of The Walking Dead?",
  INJECTION: {"head": "The Walking Dead", "type": "TV Series", "publisher": "AMC"},
  ENRICHED QUERY: "When is the last episode of season 8 of The Walking Dead? [SEP]
                   The Walking Dead  TV Series  AMC"
This family keeps expansions corpus-grounded and controllable while reducing ambiguity of short or entity-heavy queries. KGQE addresses entity ambiguity (e.g., disambiguating ”Apple” as a company vs. fruit) by injecting type-specific KG facts (e.g., ”Apple Inc. - Tech Company - Cupertino”)—a key advantage over traditional WordNet-based expansion, which fails to resolve short-query polysemy. However, KGQE struggles with low-resource domains (e.g., niche medical subfields) where KG coverage is sparse; in such cases, hybrid schemes (e.g., KGQE + BioBERT term selection [38]) are required to maintain effectiveness.
Cross-lingual KG-aware expansion (CL-KGQE)
CL-KGQE (Rahman et al., 2019) targets cross-language mismatch by combining translation, KG linking, and distributional neighbors. An input query in language $\ell_{s}$ is translated or linked cross-lingually to entities in a target KG; expansions are assembled from three sources: (i) distributional neighbors of query tokens (word embeddings), (ii) KG categories or types (e.g., DBpedia dct:subject), and (iii) hypernym or hyponym terms from lexical graphs. The final expansion pool is merged with the translated keywords and executed with a standard ranker (e.g., BM25). This hybrid design improves robustness when parallel data are scarce and terminology diverges across languages.
4.4.2. Entity-based Expansion via Hybrid KG and Document Graph
Scope and notation
In this approach, each query entity $e\in\mathcal{E}_{q}$ is first linked to a set of relevant documents or chunks containing $e$, denoted as $D(e)$. A document-level graph is then constructed where nodes correspond to these documents, and edges represent relationships such as citations, authorship, or co-occurrence of entities. Expansion terms are obtained from the graph nodes through a selection function $g(\cdot)$, and concatenated with the original query:
(42) $q^{+}=\big[q\;;\;\bigcup_{e\in\mathcal{E}_{q}}g(D(e))\big],$
where $q^{\prime}$ is the expanded query, $q;\cdot]$ denotes concatenation, and $g(\cdot)$ extracts the most relevant terms from the document graph.
KG-guided generation and retrieval (KAR)
KAR (Xia et al., 2025) leverages structural neighborhoods to guide expansion. After linking $\mathcal{E}_{q}$, retrieve an $h$-hop subgraph $\mathcal{S}_{q}=\bigcup_{e\in\mathcal{E}_{q}}\mathcal{N}_{h}(e)$ and align documents to nodes/edges. From $\mathcal{S}_{q}$, construct lightweight relational snippets (“document triples”):
(43) $T_{q}=\big\{(d_{i},r_{i,j},d_{j})\mid(v_{i},r_{i,j},v_{j})\in\mathcal{S}_{q},\ d_{i}\!\leftrightarrow\!v_{i},\ d_{j}\!\leftrightarrow\!v_{j}\big\}.$
An LLM is then prompted with $(q,T_{q})$ to write a short, knowledge-grounded expansion $g_{q}=\mathrm{LLM}(q,T_{q})$, and the final expanded query is $q^{+}=\mathrm{concat}(q;g_{q})$. Compared with pure text injection, KAR exploits relational context to surface salient, query-specific facets, which is particularly effective on semi-structured corpora (e.g., author–paper–venue graphs).
Query-specific KG construction (QSKG)
QSKG (Mackie and Dalton, 2022) builds a small, task-specific graph on the fly. A first-pass retrieval for $q$ yields $R_{k}(q)$; entities in these documents are linked and added as nodes, with edges created via within-document co-mentions or KG links, expanding iteratively to form $G_{q}$. Node salience is computed with a simple centrality or TF–IDF style score over $G_{q}$; top entities and relations are verbalized into short expansions and concatenated with the query:
(44) $q^{+}\;=\;\mathrm{concat}\Big(q;\;\big[\text{text}(v)\big]_{v\in\mathrm{TopK}(G_{q})}\Big).$
QSKG adapts to collection-specific vocabulary and senses, and tends to improve recall on ambiguous or multi-facet queries while limiting drift through graph-based salience.
4.4.3. Design considerations
KG-augmented QE improves disambiguation and coverage but introduces choices about scope and control. In practice: (i) keep injected text short (labels, types, and 1–2 discriminative facts); (ii) prefer entity linking confidence and type filters to avoid drift; (iii) in KAR/QSKG, cap graph radius and use salience to select few high-yield nodes; and (iv) for cross-lingual settings, favor KG categories and lexical graphs that transfer well across languages. These strategies preserve efficiency and mitigate noise while delivering consistent gains on entity-centric and long-tail queries.
5. Comparative Analysis: Traditional vs. PLM/LLM-Based Query Expansion
Traditional query expansion and modern PLM/LLM-driven approaches differ fundamentally in how they bridge the gap between sparse user queries and semantically rich documents. Table 2 summarizes these differences across key dimensions. Overall, traditional methods rely on static resources and lightweight heuristics, while PLM/LLM-based methods leverage parametric knowledge and contextual generation, trading efficiency for adaptability and effectiveness.
5.1. Knowledge Sources
Traditional QE depends on predefined dictionaries, ontologies, or co-occurrence statistics, which are limited in coverage and slow to update. In contrast, PLMs and LLMs internalize broad linguistic and factual knowledge through large-scale pretraining, and can be further adapted via fine-tuning or prompting, offering greater flexibility in dynamic or specialized domains.
5.2. Contextual Sensitivity
Classical methods treat terms in isolation, often failing to disambiguate polysemous words (e.g., “Apple” as fruit vs. company). PLM/LLM methods encode full-sequence context, enabling accurate disambiguation and context-aware expansions, which improves retrieval in domains with complex semantics (e.g., legal, biomedical).
Table 2. Comparison of traditional vs. PLM/LLM-based query expansion.
Dimension Traditional QE PLM/LLM-based QE
Knowledge Source Static resources, co-occurrence stats Parametric knowledge from pretraining, adaptable via prompts/fine-tuning
Contextual Sensitivity Isolated terms, context-agnostic Context-aware, sequence-level semantics
Retrieval Dependence Relies on initial ranking Reduced dependence on initial retrieval via zero-/few-shot generation
Computational Cost Lightweight, efficient Resource-intensive, higher latency
Domain Adaptability Domain-specific ontologies/dictionaries Cross-domain generalization, flexible adaptation
Retrieval Effectiveness Recall–precision trade-off Simultaneous gains in recall and precision
Integration Complexity Easy to plug into IR pipeline Requires model serving (GPUs); optionally dense indices or cross-encoders depending on pipeline
5.3. Retrieval Dependence
Pseudo-relevance feedback exemplifies traditional reliance on initial retrieval quality; poor initial results lead to topic drift. In contrast, LLMs can generate zero-shot expansions or pseudo-documents directly, reducing dependence on initial rankings and enabling more robust performance in ambiguous or long-tail queries.
5.4. Computational Cost
Traditional QE methods are lightweight and easy to deploy (e.g., RM3 with BM25). PLM/LLM-based methods, however, are resource-intensive due to large model inference and sometimes retraining, incurring higher latency and energy costs. Recent work on distillation and quantization mitigates this gap but trade-offs remain.
5.5. Domain Adaptability
Traditional QE requires curated resources for each domain (e.g., UMLS for medicine), limiting portability. PLM/LLM models, pretrained on diverse corpora, generalize better across domains and can be efficiently adapted with small amounts of domain-specific data or tailored prompts.
5.6. Retrieval Effectiveness
Traditional QE often improves recall at the expense of precision, requiring careful tuning. PLM/LLM methods, with contextual understanding and generative ability, tend to boost both recall and precision simultaneously (e.g., Query2Doc, GRF), showing stronger and more consistent gains across benchmarks.
5.7. Integration Complexity
Classical QE is plug-and-play and easily integrated into IR systems. PLM/LLM-based methods require heavier infrastructure (e.g., dense indices, model serving, GPU support), leading to higher deployment and maintenance overhead, though APIs and smaller models are reducing barriers.
6. Application Domains and Use Cases
This section reviews where query expansion matters in practice and how traditional methods and PLM/LLM-based approaches are instantiated across domains. We emphasize the mismatch patterns that QE must bridge, what evidence exists that PLM/LLM-based QE helps beyond classical techniques, and the remaining deployment constraints.
Figure 3. Application Scenarios of Query Expansion in Information Retrieval
6.1. Web Search Engines
From lexical rewriting to semantic interpretation. Commercial web search has long relied on query rewriting and synonym expansion to cope with short, diverse user queries. Early pipelines (e.g., spelling correction, query substitution) increased recall but lacked robust intent understanding, leading to noisy expansions. Modern AI-powered search integrates large neural language models (e.g., BERT, MUM) to implicitly interpret and expand queries in context (Team et al., 2023; Achiam et al., 2023), improving recall–precision balance compared to purely lexical heuristics.
Behavioral signals at scale. Large query logs provide rich signals to mine reformulations and near-synonyms from user behavior (Cui et al., 2003). Beyond pairwise mining, industrial systems aggregate candidates from behavior, taxonomy, and KBs, followed by ML-based filtering to retain expansions that improve precision. Research also leverages multi-session clicks, dwell time, and query co-occurrence to select expansion terms aligned with user intent (Kunpeng et al., 2009; Cui et al., 2003).
Event-centric expansion with PLMs for time-sensitive intents. To better handle breaking-news scenarios, Event-Centric Query Expansion (EQE) (Zhang et al., 2023) extends a query with the most salient ongoing event via a four-stage pipeline: event collection, event reformulation (leveraging pre-trained language models like BART and mT5 with prompt-tuning and contrastive learning), semantic retrieval (using a fine-tuned RoBERTa-based dual-tower model with two-stage contrastive training), and lightweight online ranking. This system has been deployed in Tencent QQ Browser Search to serve hundreds of millions of users.
6.2. Biomedical Information Retrieval
Specialized vocabulary and ontologies. Biomedical IR faces severe lay–technical vocabulary gaps (“heart attack” vs. “myocardial infarction”). Ontology-based QE (UMLS, MeSH) maps free text to controlled concepts and spelling variants (e.g., MetaMap + ATM), improving classic metrics and reducing ambiguity. Aronson and Rindflesch (1997) report up to +14.1% 11-pt AP on the Hersh collection versus non-concept baselines, illustrating the benefit of structured medical knowledge.
Domain PLMs. Domain-adapted PLMs (e.g., BioBERT, PubMedBERT) recommend contextual terms and capture fast-evolving terminology. Kelly (2021) show BioBERT-based embedding QE for rare diseases substantially improves Precision@10 (0.42 $\to$ 0.68) and nDCG (0.48 $\to$ 0.72) over BM25/MeSH, with strong gains on acronym queries. For systematic reviews, Khader et al. (2022) combine BioBERT signals with MeSH features in a learning-to-rank scheme, further boosting P@10/MAP/Recall; hybridizing semantic and ontology signals is especially effective.
Interactive LLM workflows. Ateia and Kruschwitz (2025) design an interactive biomedical RAG agent that prompts an LLM to emit an editable expanded query (Elasticsearch DSL), enabling experts to inspect and prune LLM-generated terms before BM25 retrieval. This strikes a balance between semantic priors and expert control.
Precision and hallucination control. High-stakes settings require drift control: UMLS term re-weighting with self-information can mitigate topic drift (Diao et al., 2018). For LLMs, Niu et al. (2024) propose self-refinement guided by predictive uncertainty to detect risky entities and selectively verify against medical KGs, improving factuality with limited overhead.
6.3. E-commerce Search
Catalog mismatch and normalization. User queries are short and colloquial, while product catalogs are verbose and taxonomy-driven; QE aligns lay intent with normalized attributes (Homoceanu and Balke, 2014). Synonym mining from logs plus ML filtering (e.g., eBay) improves recall and engagement in production (Mandal et al., 2019).
Generative and taxonomy-aware enrichment. Beyond synonymy, mQE-CGAN (Cakir and Gurkan, 2023) leverages GANs to synthesize expansions conditioned on lexical/semantic signals, outperforming strong baselines (e.g., BART/CTRL) on semantic similarity and coverage. ToTER (Kang et al., 2024) uses topical taxonomies for silver labeling, topic distillation, and query enrichment, improving Recall/NDCG/MAP over SPLADE++, ColBERT, PRF, and recent generative augmentation across Amazon ESCI.
LLM rewrites at scale. LLM-based personalization and on-the-fly rewrites normalize attributes and surface rare variants (Wang et al., 2024b). Incorporating intent signals (name/category prediction) with auxiliary losses improves long-tail rewriting in practice (Zhang et al., 2022). Taobao’s BEQUE (Peng et al., 2024) aligns rewrites to retrieval utility via offline feedback and Bradley–Terry-style objectives, yielding measurable GMV/CTR gains in a 14-day A/B test, with pronounced lifts on tail and few-recall queries.
6.4. Cross-Lingual & Multilingual Search
Traditional CLIR pipelines translate queries via dictionaries/MT and then apply monolingual QE; low-resource coverage and ambiguity remain pain points. In contrast, multilingual PLMs/LLMs enable zero-shot transfer and support generate-then-translate or backtranslation strategies. For example, Rajaei et al. (2024) reformulate queries via multi-language backtranslation and fuse ranked lists with RRF, improving MAP on classic testbeds (e.g., Robust04, GOV2). In practice, mixing MT with LLM-driven expansions can raise cross-language recall while controlling drift via fusion and filtering.
6.5. Open-Domain Question Answering
Answer-oriented expansions. Generation-Augmented Retrieval (GAR) (Mao et al., 2020) enriches queries with generated contexts and matches BM25-level sparse retrieval to DPR, showing that properly scoped pseudo-text can rival dense pipelines. EAR (Chuang et al., 2023) further couples expansion with reranking. Recent AQE (Yang et al., 2025) aligns expansions with downstream passage utility.
Zero-shot and process supervision. LLMs can yield high-quality expansions without task-specific fine-tuning (e.g., MILL (Jia et al., 2023)). Analyze–Generate–Refine (AGR) (Chen et al., 2024) decomposes needs and quality-controls the expansions. Hybrid text generation with PRF (HTGQE) (Zhu et al., 2023) combines multiple LLM-generated contexts (answer/sentence/title) to improve EM on NQ/Trivia.
Topic-aware ICL and vocabulary projection. TDPR (Li et al., 2025a) clusters queries to select ICL demonstrations, generates pseudo-passages, and projects embeddings onto vocabulary to produce interpretable keywords, improving R@20 and end-to-end accuracy across NQ/Trivia/WebQ/CuratedTREC.
6.6. Retrieval-Augmented Generation (RAG)
Why QE helps RAG. RAG quality hinges on retrieving the right passages; vague queries increase hallucination risk (Zhang et al., 2025a; Niu et al., 2023). Refining/expanding the query boosts retrieval coverage and answer grounding (Jagerman et al., 2023; Zhang et al., 2025b).
Representative strategies. QE-RAGpro (Hu et al., [n. d.]) uses domain-tuned embeddings to fetch Q&A exemplars for prompting LLMs to synthesize pseudo-documents, improving top-$k$ accuracy on finance datasets. QOQA (Koo et al., 2024) generates multiple LLM rewrites, selects them by query–document alignment (hybrid scores), and lifts nDCG@10 on SciFact/TREC-COVID/FiQA. KG-Infused RAG (Wu et al., 2025) performs KG-guided spreading activation, infuses KG facts into LLM expansions, and improves across multi-hop QA benchmarks. Backtranslation-based multi-view queries fused by RRF (Rajaei et al., 2024) yield consistent gains in unsupervised settings.
Practical note. Selective invocation, caching, and small/distilled models reduce latency and cost; filtering (e.g., document-aligned scoring) limits drift from generative variability.
6.7. Conversational Search
Multi-turn settings require resolving coreference, filling ellipses, and adapting to evolving intents; traditional QE struggles to selectively incorporate history. Learning to select useful turns (PRL) (Mo et al., 2023) forms expanded queries that improve MRR/NDCG and handle topic switches better. MISE (Kumar and Callan, 2020) combines Conversational Term Selection with Multi-View Reranking to fuse dialog-, passage-, and paraphrase-based views, outperforming strong neural baselines. QuReTeC (Voskarides et al., 2020) formulates term selection as lightweight classification over context and boosts MAP/NDCG@3 on CAsT.
Table 3. Representative Methods for Conversational Query Expansion
Method Granularity Model Notable strengths
PRL (Mo et al., 2023) Query BERT/ANCE Turn selection; topic-switch robustness;
higher MRR/NDCG
MISE (CTS + MVR)(Kumar and Callan, 2020) Term / Multi-view BERT/GPT-2 CTS + MVR; multi-view (dialog, passage, paraphrase);
outperforms strong neural baselines
QuReTeC(Voskarides et al., 2020) Term BERT Lightweight term selection over context;
higher MAP/NDCG@3
6.8. Code Search
From thesauri to generation. Early work enriches queries using WordNet, API names, and crowd knowledge to reduce NL–code mismatch (Lu et al., 2015; Nie et al., 2016; Lemos et al., 2014; Satter and Sakib, 2016; Zhang et al., 2017). Neural methods predict keywords and control over-expansion (Liu et al., 2019a; Huang et al., 2018, 2019).
PLM-based expansion. GACR (Li et al., 2022) augments documentation queries with generated code snippets, substantially improving CodeSearchNet across languages. SSQR (Mao et al., 2023) treats reformulation as self-supervised masked completion (T5), improving MRR on CodeBERT/Lucene without parallel data. SG-BERT/GPT2 (Liu, 2023) trains on large query corpora to generate expansions that lift top-$k$ accuracy. Hybrid embedding+sequence models (ECSQE) further enhance ranking (Bibi et al., 2025). Transformers and pretraining provide stronger semantics for bridging NL queries and code artifacts (Vaswani et al., 2017).
6.9. Lessons Learned Across Applications
(i) Traditional QE remains valuable for efficiency, transparency, and as a robust baseline (especially when latency budgets are tight or supervision is scarce). (ii) PLM/LLM-based QE offers consistent gains in contextual disambiguation, cross-domain/ cross-lingual adaptation, and recall–precision balance; interactive/inspectable workflows mitigate reliability concerns in high-stakes domains. (iii) Cost, controllability, and factuality remain the main deployment barriers. Selective invocation, fusion-based drift control, ontology/KB integration, and small/distilled models are practical levers. (iv) These observations align with our comparative analysis (Section 5) and motivate the challenges and research avenues discussed next (e.g., hallucination mitigation, feedback-grounded QE, efficiency–effectiveness trade-offs).
7. Future Directions and Open Challenges
Despite notable gains, PLM/LLM-driven QE still faces recurring weaknesses. We outline four forward-looking challenges aligned with the four dimensions of Section 4.
7.1. Reliability under unfamiliar and ambiguous queries
Evidence. In sparse or long-tail domains, LLM expansions may hallucinate entities or collapse onto a dominant sense, reducing recall and facet coverage (Abe et al., 2025). HyDE/Q2D-style pseudo-documents are often fluent but off-topic when corpus overlap is low. Directions. Retrieval-augmented prompting can inject corpus-specific evidence (Mackie et al., 2023b; Lei et al., 2024). Facet-aware diversification (multi-intent expansion, multi-query fusion) together with ambiguity detectors can trigger sense-splitting. Difficulty-aware routing may classify queries as “familiar” (prompt-only) or “unfamiliar” (feedback-grounded) (Chen et al., 2025).
7.2. Controlling expansion quality and safety
Evidence. Even strong QE can degrade precision through topic drift, while automated filtering remains unreliable (Weller et al., 2023). Directions. Lightweight judges for relevance, faithfulness, and verifiability can vet expansions pre-use. Counterfactual checks (drop-term or swap-sense) help ensure monotonic utility. In high-stakes domains, interactive editing may keep expansions faithful and auditable.
7.3. Efficiency and selective invocation
Evidence. Generative QE introduces seconds-level latency and high compute cost. Distillation shows large savings with limited loss (Pimpalkhute et al., 2024; Lee et al., 2024). Directions. Distillation and PEFT can compress LLM-QE into efficient student models. Selective invocation routes only ambiguous or tail queries to expensive QE (Chen et al., 2025), while caching and reuse exploit query-semantic clusters to amortize cost.
7.4. Dynamic domains and continual adaptation
Evidence. Static PLMs/LLMs lag behind emerging terms and events; classical PRF is corpus-specific but brittle. Directions. Continual retrieval-augmented QE blends LLM priors with on-the-fly corpus or KG evidence (Mackie et al., 2023b; Lei et al., 2024). Low-shot transfer with adapters or instruction packs supports domain-specific styles. Drift detectors can monitor expansion–click divergence and switch to conservative selection-based QE when risk is high (Naseri et al., 2021, 2022).
Takeaway.
The near-term frontier is reliable, controllable, and economical QE: corpus-grounded expansions for ambiguous queries, lightweight filters for safety, and distilled students for efficiency. Longer term, QE should adapt continually to dynamic domains while co-evolving with neural IR and RAG, guided by evaluation and governance that make expansion effects explicit.
8. Conclusion
This survey traces the evolution of query expansion from classical corpus- and lexicon-driven methods to contemporary PLM/LLM paradigms. It organized the field along four complementary dimensions: point of injection, grounding and interaction, learning and alignment, and knowledge-graph integration. These dimensions clarify how QE can be designed and combined across retrieval pipelines, from first-stage recall to re-ranking and retrieval-augmented generation.
The survey also provides a model-centric taxonomy and distilled practical guidance for deployment, including when to select explicit, implicit, or hybrid QE, how to apply feedback-grounded or constrained generation in safety-critical settings, and how to control cost through selective invocation and distillation.
Finally, the survey highlights four open challenges that define a forward agenda: improving reliability under ambiguity, strengthening quality and safety control, achieving efficiency at scale, and enabling continual adaptation to dynamic domains. Addressing these challenges will be key to advancing QE research and practice under real-world constraints.